### v1

# HARec: Hyperbolic Graph-LLM Alignment for Exploration and Exploitation in Recommender Systems

Abstract.
Modern recommendation systems often create information cocoons, limiting users’ exposure to diverse content. To enhance user experience, a crucial challenge is developing systems that can balance content exploration and exploitation, allowing users to adjust their recommendation preferences. Intuitively, this balance can be achieved through a tree-structured representation, where depth search facilitates exploitation and breadth search enables exploration. However, current works face two challenges to achieve this target: (1) Euclidean methods fail to fully capture hierarchical structures and lack flexibility in balancing exploration-exploitation, while (2) hyperbolic approaches, despite better hierarchical modeling, suffer from insufficient semantic alignment due to their reliance on Euclidean text encoders. To address these challenges, we propose HARec, a hyperbolic representation learning framework that jointly aligns user-item collaborative information with textual descriptions in hyperbolic space. Our framework introduces two key technique novelty: (1) a hierarchical-aware graph-llm alignment mechanism that enables better hierarchical representation, and (2) a hyperbolic hierarchical tree structure that facilitates user-adjustable exploration-exploitation trade-offs. Extensive experiments demonstrate that HARec consistently outperforms both Euclidean and hyperbolic baselines, achieving up to 5.49% improvement in utility metrics and 11.39% increase in diversity metrics.
Recommender Systems, Representation Learning, Hyperbolic Space
Authors affiliated with Snap Inc. served in advisory roles only for this work.
Corresponding author
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; Aug 3–Aug 7, 2025; Woodstock, NYccs: Information systems Recommender systemsccs: Information systems Collaborative filteringccs: Mathematics of computing Geometric topologyccs: Information systems Personalization
1. INTRODUCTION
As online platforms grow exponentially, users are overwhelmed by an influx of information, resulting in information cocoons that restrict their exposure to diverse and novel content (Kunaver and Požrl, 2017; Castells et al., 2021).
Traditional recommendation approaches have primarily relied on collaborative filtering and interaction data (He et al., 2020; Koren et al., 2009), where users with similar historical preferences are expected to share future interests (Chen et al., 2021). However, these methods face two critical challenges in improving user experience. First, they often fail to capture the rich semantic relationships between items and evolving user preferences. For example, a jazz enthusiast might wish to explore rock music, yet current models struggle to understand these semantic connections and adapt to such evolving preferences (as illustrated in Figure 1). Second, existing approaches inadequately capture the underlying hierarchical structure inherent in user-item networks (Zheng et al., 2017; Unger and Tuzhilin, 2020), a critical factor for improving performance and personalization.
The emergence of large language models (LLMs) offers a solution to the limitations of semantic understanding. However, integrating LLM-derived semantic features with traditional collaborative signals remains challenging, especially when capturing the hierarchical structure of user preferences and item relationships.
Hyperbolic space has emerged as a promising solution to these issues. It is a powerful tool for modeling tree-like structures and power-law distributions (Krioukov et al., 2010; Sun et al., 2021; Yang et al., 2022b, a), where its exponential volume growth naturally aligns with hierarchical, scale-free structures (Ganea et al., 2018; Chami et al., 2019; Liu et al., 2019; Zhang et al., 2021; Yang et al., 2023). In hyperbolic space, distances increase exponentially from the origin, allowing popular items or exploratory users to be positioned closer to the origin, while niche items or users with focused interests spread toward the boundaries. This spatial configuration enables hyperbolic graph neural network (GNN) models to perceive hierarchical relationships effectively, enhancing recommendation performance (Sun et al., 2021; Yang et al., 2022a, b).
However, current hyperbolic recommendation systems still face two key limitations. First, hyperbolic approaches often lack proper semantic alignment due to their dependence on Euclidean text encoders, which struggle to integrate within hyperbolic space. This mismatch leaves existing hyperbolic models vulnerable to noise in collaborative information, as they fail to capture semantic insights effectively, impacting their robustness and accuracy. Second, hyperbolic recommender systems lack a systematic approach to balancing exploration (introducing new content) and exploitation (focusing on existing preferences) (Wilson et al., 2021). This limitation hinders their adaptability to users’ evolving interests and reduces the diversity and relevance of recommended items.
Proposed Methods. To address these challenges, we propose HARec, a hyperbolic representation learning framework that facilitates exploration-exploitation through an intuitive tree-structured representation, where depth search enables exploitation and breadth search allows exploration. Our framework introduces two key components: (1) a hierarchical-aware graph-llm alignment mechanism that enhances semantic understanding and reduces noise sensitivity by jointly aligning textual descriptions with user-item collaborative information in hyperbolic space, and (2) a hyperbolic hierarchical tree structure that enables user-adjustable exploration-exploitation trade-offs. This tree structure organizes preferences hierarchically, with higher layers representing broader interests, allowing systematic navigation between exploration and exploitation.
To validate the effectiveness of HARec, we conducted extensive experiments across multiple dimensions. Our model consistently outperformed baselines in both utility and diversity, marking a significant achievement as it is the first model to excel in both aspects simultaneously. To further highlight the benefits of our hyperbolic alignment framework, we compared HARec with feature-empowered baselines, where it demonstrated superior performance. Additionally, we conducted detailed analysis on head and tail items, showing that HARec effectively boosts recommendations for tail items, thus enhancing diversity. An ablation study, along with a decoder module analysis, confirms HARec’s capability to leverage both semantic meaning and collaborative information, providing deeper insights into the advantages of our approach.
Contributions. To summarize, our primary contribution are in three aspects. The first is the proposed HARec, a hyperbolic representation learning framework that effectively balances exploration and exploitation by integrating semantic and collaborative information, supported by theoretical analysis demonstrating the advantages of this integration in hyperbolic space. In addition, we design a novel mechanism enabling personalized control over the exploration-exploitation trade-off through a hierarchical tree structure in hyperbolic space, allowing users to dynamically adjust their recommendation preferences. Moreover, through extensive experiments evaluating both accuracy and diversity metrics, we demonstrate that HARec achieves state-of-the-art performance across both criteria - achieving up to 5.49% improvement in utility metrics and 11.39% increase in diversity metrics, making it the first model to excel in both criteria simultaneously.
Figure 1. Exploration and exploitation in hyperbolic space.
2. Related Work
In this section, we review commonly used graph-based collaborative filtering approaches, large language models (LLMs) for recommendation, and hyperbolic representation learning methods.
Graph Collaborative Filtering. Graph Collaborative Filtering has emerged as a dominant approach in recommender systems, with graph neural networks (GNNs) playing a central role. Unlike traditional collaborative filtering techniques, such as matrix factorization (Koren et al., 2009), GNNs capitalize on the relational structure inherent in user-item interactions, enabling a more nuanced understanding of complex dependencies. Prominent graph-based methods, such as NGCF (Wang et al., 2019), SGL (Wu et al., 2021), and LightGCN (He et al., 2020), have been widely adopted and have consistently achieved state-of-the-art performance, significantly boosting recommendation accuracy. Nevertheless, most of these models are constrained to Euclidean space, which may fall short in capturing the underlying power-law distribution commonly seen in user-item networks.
Large Language Models for Recommendation. Large language models (LLMs) are increasingly used in recommender systems due to their strong capability in understanding textual information. Some approaches leverage LLMs directly for various downstream recommendation tasks (Bao et al., 2023; Li et al., 2023; Lyu et al., 2023; Zheng et al., 2024; Zhang et al., 2023), while others utilize LLMs as an auxiliary tool for collaborative filtering (Ren et al., 2024; Xi et al., 2024; Wei et al., 2024). However, both approaches often require substantial inference time and are computationally inefficient. In our HARec, we use language models solely to enhance data quality, avoiding the need for repeated inference during each recommendation by relying on hyperbolic embeddings for the final output.
Hyperbolic Representation Learning. To address the limitations of Euclidean space in modeling hierarchical structures, researchers have investigated the use of hyperbolic space (Chami et al., 2019; Liu et al., 2019; Yang et al., 2022b; Zhang et al., 2021), where distances grow exponentially as embeddings move away from the origin. This property makes hyperbolic geometry particularly well-suited for modeling user-item interactions by more accurately capturing latent hierarchies and dependencies. Consequently, hyperbolic representation learning has gained considerable attention in the recommender systems domain (Ying et al., 2018; Vinh Tran et al., 2020; Sun et al., 2021; Yang et al., 2022a), demonstrating improved performance over traditional Euclidean models.
However, none of these hyperbolic models fully leverage the rich semantic information embedded in language-based datasets. While some approaches incorporate knowledge graphs as supplementary information (Tai et al., 2021; Du et al., 2022), these models primarily focus on refining hierarchical structures rather than achieving a deeper semantic understanding of each item or user.
3. PRELIMINARIES
Embeddings learned from the above graph collaborative filtering are by default in the Euclidean space, which is unable to capture tree-like structures. To address the problem, we introduce how to derive these embeddings in a hyperbolic space.
Riemannian Manifold. A Riemannian manifold $(\mathcal{M},g)$ is a smooth manifold that features an inner product $g_{x}:T_{x}\mathcal{M}\times T_{x}\mathcal{M}\rightarrow\mathbb{R}$ at each point $x\in\mathcal{M}$. This inner product, known as the Riemannian metric, defines geometric characteristics of the space such as angles and curve lengths. The curvature of Riemannian manifolds results in different types of geometries: elliptic geometry with positive curvature, Euclidean geometry with zero curvature, and hyperbolic geometry with negative curvature. In this work, we focus on hyperbolic geometry, characterized by negative curvature. Hyperbolic space can be described by several equivalent models, each offering unique properties while being mathematically equivalent. The Lorentz model, also referred to as the hyperboloid model, is one of the commonly used representations of hyperbolic space.
Lorentz Model. An $n$-dimensional Lorentz manifold with negative curvature $-1/\kappa$ $(\kappa>0)$ is described as the Riemannian manifold $(\mathbb{H}^{n}_{\kappa},g_{\mathcal{L}})$. Here, $\mathbb{H}^{n}_{\kappa}=\{x\in\mathbb{R}^{n+1}:\langle x,x\rangle_{\mathcal{L}}=-\kappa,x_{0}>0\}$, $g_{\mathcal{L}}=\eta$ with $\eta=\mathbf{I}_{n}$ except $\eta_{0,0}=-1$, and $\langle\cdot,\cdot\rangle_{\mathcal{L}}$ represents the Lorentzian inner product, which is defined as by
(1) $\langle x,y\rangle_{\mathcal{L}}:=-x_{0}y_{0}+\sum_{i=1}^{n}x_{i}y_{i}.$
A tangent space is an $n$-dimensional vector space that approximates Euclidean space to hyperbolic space. Specifically, given a point $x\in\mathbb{H}_{\kappa}^{n}$, there exists a tangent space $\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}$ that approximates $\mathbb{H}_{\kappa}^{n}$:
(2) $\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}:=\{v\in\mathbb{R}^{n+1}:\langle v,x\rangle_{\mathcal{L}}=0\}.$
The tangent space is defined to be orthogonal to $x$ with respect to the Lorentzian inner product, therefore making the approximation accurate at the reference point.
We also need to define transitions between $\mathbb{H}_{\kappa}^{n}$ and $\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}$, which give rise to the exponential and logarithmic maps. For $x\in\mathbb{H}_{\kappa}^{n}$ and $v\in\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}$ with $v\neq\mathbf{0}$ and $y\neq x$, there exists a unique geodesic $\gamma:[0,1]\to\mathbb{H}_{\kappa}^{n}$ such that $\gamma(0)=x$ and $\gamma^{\prime}(0)=v$. The exponential map $\exp_{x}:\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}\to\mathbb{H}_{\kappa}^{n}$ is given by $\exp_{x}(v)=\gamma(1)$:
(3) $\exp_{x}^{\kappa}(v)=\cosh\left(\frac{\|v\|_{\mathcal{L}}}{\sqrt{\kappa}}\right)x+\sqrt{\kappa}\sinh\left(\frac{\|v\|_{\mathcal{L}}}{\sqrt{\kappa}}\right)\frac{v}{\|v\|_{\mathcal{L}}},$
where $\|v\|_{\mathcal{L}}=\sqrt{\langle v,v\rangle_{\mathcal{L}}}$ denotes the Lorentzian norm of $v$. The logarithmic map $\log_{x}$ is the inverse of the exponential map $\exp_{x}$:
(4) $\log_{x}^{\kappa}(y)=d_{\mathcal{L}}^{\kappa}(x,y)\frac{y+\frac{1}{\kappa}\langle x,y\rangle_{\mathcal{L}}x}{\|y+\frac{1}{\kappa}\langle x,y\rangle_{\mathcal{L}}x\|},$
where $d_{\mathcal{H}}^{\kappa}(x,y)$ represents the distance between $x$ and $y$ in $\mathbb{H}^{n}_{\kappa}$:
(5) $d_{\mathcal{H}}^{\kappa}(x,y)=\sqrt{\kappa}\,\text{arcosh}\left(-\langle x,y\rangle_{\mathcal{L}}/\kappa\right).$
We choose $\mathbf{o}:=\{\sqrt{\kappa},0,\ldots,0\}\in\mathbb{H}^{n}_{\kappa}$ as the reference point for all the operations, making the transitions reproducible. In this work, we set $\kappa=1$, resulting in a curvature of $-1$. Therefore, in the following discussions, we would omit $\kappa$ for simplicity.
4. METHODOLOGY
In this section, we present a comprehensive overview of our proposed model, HARec. Our approach not only aims to deliver accurate recommendations but also addresses users’ growing demand for more diverse suggestions. Furthermore, by constructing a hierarchical tree from the learned embeddings, we introduce an additional mechanism that enables users to control the degree of exploration in the recommended items. The overall framework is in Figure 2.
Figure 2. The overall architecture of HARec. (i) Hyperbolic Graph Collaborative Filtering: Encodes collaborative information using hyperbolic GNNs; (ii) Hyperbolic Alignment: Aligns semantic and collaborative information within hyperbolic space; (iii) Hyperbolic Hierarchical Structure: Builds hierarchy tree from hyperbolic embeddings.
4.1. Hyperbolic Graph Collaborative Filtering
Similar to Euclidean space, hyperbolic graph collaborative filtering captures high-order dependencies between users and items via a message aggregation mechanism. In our method, we use the Lorentz representation for both users and items.
Hyperbolic Embedding Initialization. As discussed in section 3, we fix the origin $\mathbf{o}:=\{\sqrt{\kappa},0,\ldots,0\}\in\mathbb{H}^{n}_{\kappa}$ and use it as the reference point. We adopt the Gaussian sampling method for the initialization of embeddings in Euclidean space, after that, by injecting them into the tangent space of the reference point, we derive the initial hyperbolic embeddings of both users and items:
(6) $z_{u}^{0}=(0,e_{u}),\quad\qquad z_{i}^{0}$ $=(0,e_{i})$
$h_{u}^{0}=\exp_{\mathbf{o}}(z_{u}^{0}),\quad\quad h_{i}^{0}$ $=\exp_{\mathbf{o}}(z_{i}^{0})$
where $e_{u},e_{i}$ are Gaussian distributed initialized Euclidean embeddings, and $z_{u}^{0},z_{i}^{0}$ are corresponding embeddings in the tangent space of the origin point. $h_{u}^{0},h_{i}^{0}$ therefore stands for the resulting initialized hyperbolic embeddings.
Hyperbolic Message Passing. Message passing between each node is realized through hyperbolic neighborhood aggregation, which is used to extract explicit user-item interaction. Given user $u$, item $i$, as well as their neighborhood $\mathcal{N}_{u},\mathcal{N}_{i}$, their embeddings at layer $l$ in the tangent space are updated as follows:
(7) $z_{u}^{l}=z_{u}^{l-1}+\sum_{i\in\mathcal{N}_{u}}\frac{1}{|\mathcal{N}_{u}|}z_{i}^{l-1},\quad\quad z_{i}^{l}=z_{i}^{l-1}+\sum_{u\in\mathcal{N}_{i}}\frac{1}{|\mathcal{N}_{i}|}z_{u}^{l-1}$
where $|\mathcal{N}_{u}|$ and $|\mathcal{N}_{i}|$ represent the number of neighbors of $u$ and $i$, respectively. The final tangent space embedding is derived by summing up embeddings at each layer:
(8) $z_{u}=\sum_{l}z_{u}^{l},\quad\quad z_{i}=\sum_{l}z_{i}^{l}$
Since $z_{u}$ and $z_{i}$ are on the tangent space, we still need to project it into hyperbolic space to derive the hyperbolic embeddings:
(9) $h_{u}=\exp_{\mathbf{o}}(z_{u}),\quad\quad h_{i}=\exp_{\mathbf{o}}(z_{i})$
Hyperbolic Prediction Function. Through hyperbolic messaging passing, we’ve captured high-order relational information into hyperbolic embeddings. Hyperbolic distance is now used to define a prediction function, which is used to predict the possibility that a user would interact with an item:
(10) $p(u,i)=\frac{1}{d_{\mathcal{H}}(h_{u},h_{i})}$
Hyperbolic Margin Ranking Loss. Bayesian Personalized Ranking (BPR) loss has been widely applied in recommender systems (Rendle et al., 2012), aiming to pull together user-item pairs with recorded interactions while pushing apart those without. In our model, we adopt the hyperbolic margin ranking loss (Sun et al., 2021), designed to prevent the undesirable collapse seen in Euclidean space by adding an additional margin to the original BPR loss:
(11) $\ell_{margin}(u,i,j)=\max(d_{\mathcal{H}}^{2}(h_{u},h_{i})-d_{\mathcal{H}}^{2}(h_{u},h_{j})+m,0),$
where $d_{\mathcal{H}}$ represents hyperbolic distance, $m$ is the margin, $(u,i)$ is an observed interaction, and $(u,j)$ is a sampled negative pair.
Hyperbolic Informative Mechanism. We further integrate a geometric-aware mechanism (Yang et al., 2022a) to enhance the pull-push process, consisting of two parts: Hyperbolic-aware Margin Learning (HAML) and Hyperbolic Informative Negative Sampling (HINS). Given that distances in hyperbolic space grow exponentially as embeddings deviate from the origin, we aim to assign a larger margin when user-item pairs are near the origin. HAML addresses this issue by dynamically adapting the margin as follows:
(12) $m^{\mathcal{H}}_{ui}$ $=\mathrm{sigmoid}(\delta),$
$\delta$ $=\frac{d_{\mathcal{H}}^{2}(e_{u},\mathbf{o})+d_{\mathcal{H}}^{2}(e_{i},\mathbf{o})-d_{\mathcal{H}}^{2}(e_{u},e_{i})}{e_{u,0}\cdot e_{i,0}}$
where $e_{u,0}$ and $e_{i,0}$ represent the zeroth coordinate elements of $e_{u}$ and $e_{i}$, respectively, corresponding to the embeddings norms in the hyperbolic space. HINS is an algorithm that samples negative nodes close to the positive item, yielding informative negative samples for both head and tail items. This is achieved by iteratively calculating the distances between the positive item and all candidate negative samples until the nearest one is identified.
4.2. Semantic Representations Generation
In this subsection, we introduce a paradigm for distilling useful information from raw text descriptions of items and users into dense semantic embeddings, which will be used later in our alignment framework that derives the final embeddings, integrating both collaborative information and semantic insights. Inspired by prior research (Ma et al., 2024; Ren et al., 2024) that leverages the power of large language models (LLMs) to extract useful user/item descriptions from raw, noisy textual data, we adopt a similar framework to generate user/item profiles with rich textual descriptions:
(13) $\mathcal{P}_{i}=\textit{LLMs}(\mathcal{S}_{i},\mathcal{T}_{i})$
where $\mathcal{S}_{i}$ represents system prompts to the LLMs, $\mathcal{T}_{i}$ denotes raw textual information, and $\mathcal{P}_{i}$ is the resulting user/item profile, consisting of sentence-level descriptions.
While we’ve derived textual information for users/items, it cannot be directly fused with the embeddings generated from hyperbolic collaborative filtering. To integrate both types of knowledge, we encode the textual information into embeddings using pretrained text encoder models, such as BERT (Devlin, 2018):
(14) $e_{i}=\textit{Encoder}(\mathcal{P}_{i})$
4.3. Hyperbolic Alignment
Given two sets of embeddings: hyperbolic embeddings with collaborative information and semantic embeddings with textual information, we aim to maximize the mutual information between them, creating unified embeddings that integrate both types of knowledge. To achieve this, we introduce a hyperbolic alignment loss, which mitigates the effects of noisy signals in representation learning and improves recommendations for tail items by aligning information within the same hyperbolic space. A theoretical proof of the advantages of alignment in hyperbolic space over Euclidean space is provided in the next subsection.
Before calculating the alignment loss, two steps are necessary: 1) the semantic embeddings must be adjusted to match the dimensionality of the collaborative embeddings, and 2) they should be transformed into a common representation space. To accomplish this, we first apply a multi-layer perceptron (MLP) to the semantic embeddings to ensure dimensional alignment. We then project the semantic embeddings into hyperbolic space so they can align with the hyperbolic collaborative embeddings:
(15) $s_{i}$ $=\exp_{x}(e_{i})$
$s^{\prime}_{i}$ $=\textit{MLP}(s_{i})$
where $\exp_{x}$ is defined as in 3, $e_{i}$ represents the semantic embeddings from the encoder, $s_{i}$ is the dimension-aligned semantic embedding, and $s^{\prime}_{i}$ denotes the final projected semantic embedding in hyperbolic space. The hyperbolic alignment loss is then calculated as:
(16) $\ell_{align}(i)=d_{\mathcal{H}}^{2}(h_{i},\textbf{s}^{\prime}_{i})$
where $d_{\mathcal{H}}$ represents the hyperbolic distance, and $h_{i}$ is the hyperbolic embedding with collaborative information.
4.4. Gradient Analysis of Semantic Alignment.
To further explore the impact of aligning semantic embeddings within the hyperbolic model, we analyze the gradient behavior with respect to nodes of varying norms in Proposition 4.1. This provides insight into the hierarchical preservation properties of the hyperbolic space, which are absent in Euclidean space.
Proposition 4.1.
The gradient magnitude of semantic alignment in hyperbolic space $\|\nabla_{\mathbf{x}}d_{\mathcal{H}}\|\approx\frac{\|\hat{\mathbf{x}}-\hat{\mathbf{y}}\|}{\|\mathbf{x}\|(1-\cos\theta)}$ enables adaptive updates for hierarchical preservation, while Euclidean space maintains constant magnitude $\|\nabla_{\mathbf{x}}d_{E}\|=1$, where the $d_{\mathcal{H}}$ and $d_{\mathcal{E}}$ represent the hyperbolic distance and Euclidean distance, respectively.
Hyperbolic Space. Suppose $\mathbf{x}$ is a hyperbolic embedding derived from  4.1, and $\mathbf{y}$ is the semantic embedding from  4.2. The hyperbolic distance is given by:
(17) $d_{\mathcal{H}}(\mathbf{x},\mathbf{y})=\operatorname{arccosh}(z)$
where $z=-\langle\mathbf{x},\mathbf{y}\rangle_{\mathcal{L}}=x_{0}y_{0}-\sum_{i=1}^{n}x_{i}y_{i}$. For large $\|\mathbf{x}\|$ and $\|\mathbf{y}\|$, given $x_{0}=\sqrt{1+\|\mathbf{x}\|^{2}}$ and $y_{0}=\sqrt{1+\|\mathbf{y}\|^{2}}$, we adopt the following approximations for further calculation:
(18) $x_{0}\approx\|\mathbf{x}\|,\quad y_{0}\approx\|\mathbf{y}\|,\quad\sqrt{z^{2}-1}\approx z$
(19) $z=x_{0}y_{0}-\sum_{i=1}^{n}x_{i}y_{i}\approx\|\mathbf{x}\|\|\mathbf{y}\|(1-\cos\theta)$
Let $x_{j}=\|\mathbf{x}\|\hat{x}_{j}$ and $y_{j}=\|\mathbf{y}\|\hat{y}_{j}$ (where $\hat{x}_{j}$ represents the direction of $x_{j}$). The partial derivative of $z$ with respect to $x_{j}$ is:
(20) $\frac{\partial z}{\partial x_{j}}=\frac{x_{j}}{x_{0}}y_{0}-y_{j}\approx\frac{x_{j}}{\|\mathbf{x}\|}\|\mathbf{y}\|-y_{j}=\|\mathbf{y}\|(\hat{x}_{j}-\hat{y}_{j})$
Thus, the gradient norm of $d_{\mathcal{H}}$ for $\mathbf{x}$ can be approximated by:
(21) $\|\nabla_{\mathbf{x}}d_{\mathcal{H}}\|=\frac{\|\nabla_{\mathbf{x}}z\|}{\sqrt{z^{2}-1}}\approx\frac{\|\nabla_{\mathbf{x}}z\|}{z}=\frac{\|\mathbf{y}\|\|\hat{\mathbf{x}}-\hat{\mathbf{y}}\|}{\|\mathbf{x}\|\|\mathbf{y}\|(1-\cos\theta)}=\frac{\|\hat{\mathbf{x}}-\hat{\mathbf{y}}\|}{\|\mathbf{x}\|(1-\cos\theta)}$
This result indicates that nodes with large norms undergo smaller gradient updates, preserving local structures, while nodes with smaller norms adjust more, refining global hierarchical relationships and capturing hierarchical structure effectively.
Euclidean Space. For comparison, we examine the gradient behavior in Euclidean space. The Euclidean distance is defined as:
(22) $f(\mathbf{x})=d_{E}(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|$
The gradient of $f(\mathbf{x})$ with respect to $\mathbf{x}$ is:
(23) $\nabla_{\mathbf{x}}f(\mathbf{x})=\frac{\mathbf{x}-\mathbf{y}}{\|\mathbf{x}-\mathbf{y}\|}=\frac{\mathbf{x}-\mathbf{y}}{d_{E}(\mathbf{x},\mathbf{y})}$
which has a magnitude of:
(24) $\|\nabla_{\mathbf{x}}f(\mathbf{x})\|=1$
In summary, the analysis shows that hyperbolic space provides adaptive gradient magnitudes that depend on the norms of the nodes, effectively capturing and preserving hierarchical structures. In contrast, Euclidean space, with its constant gradient magnitude, cannot adapt to hierarchical nuances, making it less suitable for modeling hierarchical data in recommender systems.
4.5. Hyperbolic Hierarchical Structure
Figure 3. Hierarchical Structure Visualization.
In this subsection, we introduce a mechanism that allows users to control the level of exploration in their recommendations by constructing a hierarchy tree based on the learned hyperbolic embeddings, an example is shown in Figure 3. These embeddings integrate both collaborative information and language semantics, enabling more nuanced recommendations.
Hierarchy Tree. The primary motivation for constructing a hierarchy tree is to uncover hidden structures in the data that are not explicitly reflected in the dataset. For instance, consider the Amazon-CD dataset: a CD might be categorized under genres like “folk music”, but it cannot be labeled as “folk music preferred by teenagers” or further refined to “folk music preferred by teenagers who live in large cities”. However, in reality, such nuanced preferences exist within the collaborative information. When recommending based solely on past interactions, we may overlook these hidden layers of structure that are implicit in the dataset (e.g., “music” → “folk music” → “folk music preferred by teenagers” → “folk music preferred by teenagers who live in large cities”).
As previously discussed, in hyperbolic space, the norm of the embeddings reflects the level of exploration, where depth search facilitates exploitation and breadth search enables exploration. Building on this concept, we use the learned embeddings to form the bottom layer of the hierarchy tree. We then apply K-Means clustering in hyperbolic space to derive higher-level cluster nodes, which act as pseudo clustering nodes within the hierarchy, representing group preferences. For example, if an item is classified as “folk music preferred by teenagers who live in large cities”, its parent cluster may represent “folk music preferred by teenagers”, while its sibling clusters could include “folk music preferred by teenagers who live in small towns”. This hierarchical clustering process continues iteratively until reaching the root, thereby creating a multi-level tree structure that captures different levels of exploration.
Exploration-Exploitation Balance. To allow users control over the balance of exploration and exploitation, we introduce two adjustable parameters: temperature $\tau$ and hierarchy level $l$. In the standard recommendation approach, items are recommended based on similarity to the user’s embedding, which may limit exploration. For users with focused preferences, this method may consistently recommend items that align too closely with their past choices, offering little opportunity to explore different categories.
To address this issue, we modify the recommendation process by replacing a portion of the recommended items with alternatives from other branches of the hierarchy tree. This modification gives users control over the balance between exploration and exploitation. Temperature $\tau$ allows users to specify the proportion of original recommendations to replace, while level $l$ determines the extent of exploration within the hierarchy. For example, if a user selects $\tau=0.5$ and $l=2$, half of the originally recommended items will be retained, while the other half will be replaced by items sampled from nodes two levels up in the hierarchy tree, with selections made from the connected bottom-layer nodes.
5. EXPERIMENTS
Table 1. Comprehensive model comparison based on utility and diversity metrics, where higher values signify superior performance. The best-performing results for each metric are highlighted in bold, with the second-best results underlined.
Data Models utility $\uparrow$ Diveristy $\uparrow$
Recall@10 NDCG@10 Recall@20 NDCG@20 Div@10 H@10 EPC@10 Div@20 H@20 EPC@20
Amazon- books LightGCN 0.0950 0.0726 0.1428 0.0884 0.4469 9.8554 0.7429 0.4504 10.485 0.7798
TAG-CF 0.0986 0.0755 0.1468 0.0912 0.4090 10.915 0.7899 0.4517 11.465 0.8227
LightGCL 0.1075 0.0846 0.1515 0.0991 0.4651 11.511 0.8523 0.4974 11.941 0.8728
SimGCL 0.0997 0.0762 0.1478 0.0918 0.4286 10.341 0.7757 0.4802 10.915 0.8076
HCCF 0.1045 0.0813 0.1488 0.0962 0.5008 11.814 0.8579 0.5433 12.137 0.8746
HGCF 0.1013 0.0759 0.1511 0.0921 0.4742 11.577 0.8623 0.5041 11.732 0.8585
HICF 0.1113 0.0862 0.1599 0.1022 0.4702 11.922 0.8561 0.5261 12.129 0.8660
Ours 0.1167 0.0902 0.1682 0.1069 0.5255 12.096 0.8700 0.5754 12.254 0.8758
Yelp LightGCN 0.0678 0.0565 0.1135 0.0717 0.2603 8.3263 0.6842 0.2277 9.1811 0.7343
TAG-CF 0.0692 0.0571 0.1142 0.0722 0.2353 8.4516 0.6939 0.2276 9.3108 0.7428
LightGCL 0.0709 0.0594 0.1141 0.0739 0.2856 11.243 0.8281 0.3064 11.823 0.8550
SimGCL 0.0707 0.0586 0.1184 0.0748 0.3397 9.8191 0.7642 0.3460 10.508 0.7973
HCCF 0.0680 0.0562 0.1097 0.0701 0.3168 11.942 0.8572 0.3535 12.311 0.8752
HGCF 0.0654 0.0510 0.1099 0.0665 0.3133 11.977 0.8718 0.3289 12.103 0.8747
HICF 0.0709 0.0579 0.1169 0.0737 0.3478 11.877 0.8557 0.3721 12.147 0.8666
Ours 0.0739 0.0597 0.1224 0.0762 0.3874 12.155 0.8716 0.3961 12.329 0.8773
Google- reviews LightGCN 0.0932 0.0745 0.1430 0.0929 0.2422 9.9149 0.7241 0.2725 10.761 0.7749
TAG-CF 0.0942 0.0754 0.1447 0.0940 0.2376 10.184 0.7485 0.2497 11.031 0.7961
LightGCL 0.0945 0.0768 0.1461 0.0959 0.2639 11.643 0.8333 0.2705 12.317 0.8655
SimGCL 0.0979 0.0784 0.1482 0.0971 0.2966 10.361 0.7486 0.2950 11.119 0.7923
HCCF 0.0923 0.0744 0.1446 0.0937 0.2195 10.741 0.8056 0.2402 11.561 0.8433
HGCF 0.0910 0.0697 0.1485 0.0909 0.2859 12.248 0.8538 0.3321 12.474 0.8619
HICF 0.0951 0.0736 0.1515 0.0944 0.3262 12.376 0.8523 0.3510 12.660 0.8640
Ours 0.1002 0.0772 0.1598 0.0992 0.3185 12.466 0.8598 0.3543 12.769 0.8728
Table 2. Comparison with feature-enhanced Euclidean baselines, with the best-performing results highlighted in bold.
Data Amazon-books Yelp Google-reviews
Metrics R@20 N@20 Div@20 H@20 EPC@20 R@20 N@20 Div@20 H@20 EPC@20 R@20 N@20 Div@20 H@20 EPC@20
LightGCN+ 0.1466 0.0901 0.4349 10.3915 0.7802 0.1156 0.0733 0.2697 9.6903 0.7526 0.1438 0.0940 0.2707 10.8664 0.7801
LightGCL+ 0.1545 0.0992 0.5064 11.9761 0.8728 0.1131 0.0739 0.2759 11.8617 0.8581 0.1443 0.0946 0.2744 12.4700 0.8776
SimGCL+ 0.1494 0.0926 0.5163 11.4751 0.8342 0.1195 0.0752 0.3152 9.9952 0.7811 0.1504 0.0982 0.3030 11.2111 0.7953
HCCF+ 0.1525 0.0940 0.4316 11.7984 0.8616 0.1170 0.0729 0.2884 11.6906 0.8577 0.1438 0.0941 0.2524 11.5433 0.8406
Ours 0.1682 0.1069 0.5754 12.2543 0.8758 0.1224 0.0762 0.3961 12.3290 0.8773 0.1598 0.0992 0.3543 12.7690 0.8728
5.1. Experimental Settings
Datasets. To evaluate our proposed model, we utilize three widely-used public datasets: Amazon-books, which contains user purchase behaviors within the book category on Amazon; Yelp, which records customer ratings and reviews for restaurants on Yelp; and Google-reviews, which includes user reviews and business metadata from Google Maps. Detailed statistics for these datasets , including the distribution of head and tail items, are provided in Table 3. The high proportion of tail users (T80) across these datasets highlights the prevalence of long-tail distributions, supporting the assumption of a power-law distribution in real-world data.
Evaluation Metrics. We assess our recommended items using both utility and diversity metrics, demonstrating that our model achieves precise recommendations for users’ preference while enhancing diversity in the recommendation list.
For utility, we adopt two widely-used ranking metrics: Recall and NDCG, which measure model effectiveness. To evaluate diversity, we employ three metrics: Distance Diversity (Ziegler et al., 2005; Zhang and Hurley, 2008), Shannon Entropy (Mehta et al., 2012), and Expected Popularity Complement (Vargas and Castells, 2011).
Distance Diversity (Div) quantifies the variation among recommended items, encouraging diversity within the recommendation:
(25) $Div(u)=\frac{\sum_{i\in R_{u}}\sum_{j\in R_{u},j\neq i}d(i,j)}{\text{Number of}(i,j)\text{ pairs}}$
where $R_{u}$ is the recommendation list for user $u$, and $d(i,j)$ represents the distance between items $i$ and $j$. Number of $(i,j)$ pairs is calculated as $\binom{|R_{u}|}{2}=\frac{|R_{u}|(|R_{u}|-1)}{2}$. For fair comparison, we compute distances within the same representation space across models.
Shannon Entropy (H) quantifies the unpredictability within the recommendations, with higher entropy signifying greater diversity:
(26) $H=-\sum_{i\in\text{set}(\sum_{u}R_{u})}p(i)\log p(i)$
where $\sum_{u}R_{u}$ denotes all recommended items across users (including duplicates), and $p(i)=\text{count}(i)/\sum_{i\in\text{set}(\sum_{u}R_{u})}\text{count}(i)$ is the proportion of item $i$ among all recommended items.
Expected Popularity Complement (EPC) assesses the extent of bias towards popular items by promoting niche recommendations:
(27) $EPC=1-\frac{1}{|\text{set}(\sum R_{u})|}\sum_{i\in\text{set}(\sum_{u}R_{u})}\frac{\text{pop}(i)}{\max_{i}(\text{pop}(i))}$
where $N_{u}$ is the set of items user $u$ has interacted with, and $\text{pop}(i)=\text{count}(i)/\sum_{i\in\text{set}(\sum_{u}N_{u})}\text{count}(i)$ denotes item $i$’s popularity.
Baselines. We compare our model with state-of-the-art baselines in Euclidean and Hyperbolic spaces. Euclidean models include:
  • LightGCN (He et al., 2020): A lightweight graph recommender model by removing redundant neural modules in graph convolution.
  • TAG-CF (Ju et al., 2024): Enhances collaborative filtering by performing a single message-passing step only at inference time.
  • LightGCL (Cai et al., 2023): Uses singular value decomposition (SVD) for robust graph augmentation without data augmentations.
  • SimGCL (Yu et al., 2022): Replaces graph augmentations in contrastive learning by introducing random noise during initialization.
  • HCCF (Xia et al., 2022): Integrates hypergraph and contrastive learning, capturing both local and global user dependencies.
Hyperbolic GNN models include:
  • HGCF (Sun et al., 2021): Combines hypergraph learning with collaborative filtering using hyperbolic margin ranking loss.
  • HICF (Yang et al., 2022a): Enhances head and tail item performance with hyperbolic geometry, improving utility for long-tail items.
Implementation Details. We set embedding dimensions to 50 and a training batch size of 1024. Embeddings in hyperbolic space are optimized using Riemannian SGD (RSGD) (Bécigneul and Ganea, 2018; Sun et al., 2021), while the MLP adapter is optimized with Adam (Kingma, 2014). Training incorporates early stopping technique based on Recall@10 with 10 patience steps. For the hierarchical structure, we use a $1:8$ ratio for parent and child nodes. We generate user and item profiles with GPT-3.5-turbo and encode them with text-embedding-ada-002 (Neelakantan et al., 2022) for Amazon-books and Yelp, and BERT (Devlin, 2018) for Google-reviews. The profile decoder backbone is Llama 3.1-8B. All experiments are conducted on NVIDIA A40/A100 GPUs with 40GB memory.
Table 3. Statistics of the experimental data.
Dataset #User #Item #Interactions Density
All H20(%) T80(%)
[…]
5.2. Performance Comparison
Overall Comparison. The performance comparison across all models is presented in Table 1. Our HARec consistently outperforms the baselines in both utility and diversity metrics, achieving up to $5.49\%$ improvement in utility metrics and $11.39\%$ increase in diversity metrics. Notably, most baselines struggle to achieve strong performance in both utility and diversity simultaneously. For instance, while HICF performs second-best in utility metrics on the Amazon-books dataset, outperforming other baselines by a significant margin (excluding our model), it lacks the same advantage in diversity metrics, where several baselines achieve better results. This highlights the superiority of our model, which achieves state-of-the-art performance in both utility and diversity: a feat not previously attained by any baseline.
Baselines with Feature Enhancement. To ensure a fair comparison, we further enhance Euclidean baselines by integrating feature information. Specifically, we align semantic embeddings with the original Euclidean embeddings following the structure in Section 4.2, but in Euclidean space. The results are presented in Table 2, with enhanced baselines denoted by “+”. The data shows that our HARec still surpasses all baselines in both utility and diversity. Interestingly, while the feature-enhanced versions of the Euclidean baselines generally improve in utility metrics, they do not exhibit a stable improvement in diversity metrics. This aligns with our theoretical analysis in Section 4.5, which demonstrates the advantages of our alignment model in the hyperbolic space.
Table 4. Layer-wise norms within the hierarchy structure.
Data level-0 level-1 level-2 level-3 level-4
Amazon-books 0.2388 0.4859 0.5143 0.5889 0.6773
Yelp 0.1758 0.3500 0.4719 0.5590 0.6307
Google-reviews 0.4374 0.8585 1.0682 1.1731 1.2479
Figure 4. Ablation study on model variants.
Table 5. Performance on H20 (head) and T80 (tail) items, with the top hyperbolic results highlighted in bold. $\Delta_{\mathcal{H}}(\%)$ represents the improvement over the second-best hyperbolic baseline, which is underlined.
Data Amazon-books Yelp Google-reviews
Metrics Recall@20 NDCG@20 Recall@20 NDCG@20 Recall@20 NDCG@20
H20 T80 H20 T80 H20 T80 H20 T80 H20 T80 H20 T80
LightGCN 0.1162 0.0266 0.0744 0.0140 0.1076 0.0059 0.0693 0.0024 0.1311 0.0119 0.0878 0.0051
HGCF 0.1083 0.0428 0.0673 0.0248 0.0789 0.0310 0.0489 0.0176 0.1123 0.0362 0.0714 0.0195
HICF 0.1086 0.0513 0.0713 0.0309 0.0847 0.0322 0.0556 0.0180 0.1147 0.0388 0.0746 0.0208
Ours 0.1125 0.0557 0.0730 0.0339 0.0863 0.0361 0.0558 0.0204 0.1155 0.0443 0.0755 0.0238
$\Delta_{\mathcal{H}}(\%)$ +3.59 +8.58 +2.38 +9.71 +1.89 +12.11 +0.36 +13.33 +0.70 +14.18 +1.21 +14.42
Figure 5. Performance across different levels of exploration.
5.3. Ablation Study
We perform an ablation study to examine the contributions of two key components in our model: hyperbolic margin ranking loss and semantic alignment loss. To maintain meaningful evaluation, we focus solely on utility metrics, as random recommendations may artificially inflate diversity without practical utility. The following notations are used for our ablation variants: “w/o M” denotes the model without hyperbolic margin ranking loss, “w/o S” represents the model without semantic alignment loss, and “w/o M & S” indicates that both margin ranking and semantic alignment are excluded, leaving only the BPR loss.
Figure 4 shows that our full model outperforms all ablation variants, highlighting its overall effectiveness. Models lacking both margin ranking and semantic alignment perform the worst, while removing either module results in decreased utility, emphasizing the importance of each component and their effective integration.
5.4. Hierarchical Structure Analysis
To validate the effectiveness of our hierarchical structure, which allows users to control their level of exploration within recommendations, we conducted a series of analyses. Our goal is to demonstrate how the hierarchical model provides a structured balance between exploration and exploitation, allowing users to adjust the generality of recommendations by themselves.
First, we analyze the norms of node embeddings across hierarchical layers. As shown in Table 4, nodes in higher layers have consistently smaller norms. This observation supports our design intention: higher-layer nodes represent broader, generalized group preferences, aligning with findings in Section 4.5. This structure enables more abstract grouping as the layers ascend, thereby capturing progressively broader user interests.
To further examine the hierarchy’s practical utility, we designed experiments that simulate varying user preferences in daily interactions. Using a temperature parameter $\tau=0.2$ to control recommendation randomness, we tested recommendation performance across different hierarchical levels, from layer 00 (original recommendation level) up to layer $3$. The results, visualized in Figure 5, reveal a clear pattern: as the level $l$ increases, utility metrics tend to decrease, while diversity metrics increase.
This progression highlights the intended trade-off our model offers: recommendations from lower layers (e.g., layer 00) focus on utility and specific user interests, ideal for exploitation. In contrast, higher layers (e.g., layer $3$) prioritize diversity by presenting broader, less tailored options, thus encouraging exploration. This systematic change across levels provides users with a flexible mechanism to adjust recommendation focus: from niche, highly relevant items to a more varied, exploratory set—based on their current preference for exploration or exploitation.
To underscore the effectiveness of semantic alignment, particularly for cold-start (tail) items, we conducted experiments measuring utility performance on both head and tail items. In our context, head items (H20) refer to the top $20\%$ of items with the highest number of interactions in the training dataset, representing the most popular items. The remaining items are categorized as tail items (T80), which have limited interaction history, representing less frequently engaged (cold-start) items.
For simplicity, we compare our model with all hyperbolic baselines and a representative Euclidean baseline (LightGCN). The results, displayed in Table 5, reveal significant improvements, especially for tail items. Compared with hyperbolic baselines without semantic alignment, our model shows a notably larger performance boost on tail items than on head items. This finding indicates that semantic alignment particularly enhances the recommendation of cold-start items, effectively mitigating the cold-start problem for items with limited or no prior interactions.
6. CONCLUSION
In this work, we introduce HARec, a novel hyperbolic graph-text alignment framework that effectively balances exploration and exploitation by integrating semantic and collaborative information, with theoretical analysis supporting the advantages of this approach within hyperbolic space. Additionally, we propose a unique mechanism that enables personalized control over the exploration-exploitation trade-off through a hierarchical tree structure, allowing users to adjust recommendation preferences dynamically. Extensive experiments demonstrate that HARec achieves state-of-the-art performance, with up to a $5.49\%$ improvement in utility metrics and an $11.39\%$ increase in diversity metrics, making it the first model to excel in both criteria simultaneously, a significant achievement that has yet to be attained by any previous models.
Appendix A Appendix
In the supplementary materials, we provide additional details on our datasets and present several case studies illustrating the performance of our decoder module.
A.1. Profile Decoder
To further evaluate the expressiveness of our learned hyperbolic embeddings, we introduce a profile decoder designed to reconstruct the original profile from these embeddings, thereby broadening their application beyond recommendation tasks. Our method begins by mapping the hyperbolic embeddings back into Euclidean space to facilitate alignment with traditional network layers. These Euclidean-mapped embeddings are then processed through a mixture of experts (MoE) module. The MoE serves to refine the embeddings, ensuring consistency in both dimensionality and semantic representation (Ma et al., 2024). Once processed, these embeddings are fed into the language model, specifically after the positional layer, to incorporate them within a sequential architecture. To enhance training efficiency and minimize computational cost, we freeze the language model’s parameters, making the MoE module the only trainable component in the architecture.
Details of the input prompt structure and processing flow are depicted in Figure 6. To facilitate seamless embedding integration, we introduce two additional tokens, ¡USER_EMBED¿ and ¡ITEM_EMBED¿, to the language model’s tokenizer. This addition enables these embeddings to be treated as single, coherent tokens, ensuring accurate tokenization and embedding interpretation. After passing through the positional layer, the placeholders are dynamically replaced by the adapted embeddings generated by the MoE, allowing the embeddings to retain context-specific information.
A.2. Decoder Quality
To assess the effectiveness of our decoder mechanism, we compare the embeddings from our model against those from Euclidean embeddings (LightGCN) and Hyperbolic embeddings (HICF) as inputs to the decoder. The results are shown in Table 6. We use BERTScore (Zhang et al., [n. d.]) to evaluate the semantic alignment of generated sentences with the original profile. Results indicate that our HARec outperforms both Euclidean and hyperbolic baselines, demonstrating that our model’s embeddings support recommendation and retain user/item profile information. This finding underscores that our model captures collaborative information alongside recoverable language semantics: a capability not achieved by previous models.
A.2.1. Decoder Case Study
To vividly illustrate the semantic capacity of our embeddings, we present two case studies demonstrating the model’s ability to recover user and item profiles.
————————————————————————————————————— Original user profile:
Based on the user’s interactions and reviews, this user is likely to enjoy unique dining experiences with a variety of options, trendy atmosphere with good drink and food selections, premium quality frozen yogurt with a wide variety of toppings, creative burgers and hot dogs, delicious Thai cuisine, and authentic and spicy Chinese cuisine served family-style.
Recovered user profile:
Based on the user’s interactions, it is likely that they would enjoy businesses that offer a variety of cuisines, including Mexican, Italian, and Asian fusion. They appreciate good food, drinks, and a relaxed atmosphere. They also value quality and freshness in their food.
————————————————————————————————————— In this example, the profile recovered by our decoder accurately captures the user’s preferences for Thai and Chinese cuisines, broadly summarized as an affinity for Asian fusion flavours. Additionally, our decoder identifies the user’s interest in diverse culinary experiences, indicating a propensity toward exploration.
————————————————————————————————————— Original item profile:
Ceol Irish Pub in Reno is a popular bar and pub known for its happy hour deals, great service, and friendly atmosphere. It is suitable for those who enjoy trivia nights, live music, and a wide selection of drinks, particularly whiskey. The pub has a welcoming environment, with bartenders who are attentive and quick to serve.
Recovered item profile:
Users who enjoy a variety of beer options, live music, and a relaxed atmosphere would enjoy Ceol Irish Pub.
————————————————————————————————————— The alignment between the original and recovered item profiles shows that our decoder effectively captures key thematic elements of Ceol Irish Pub. The original profile emphasizes aspects such as happy hour, a wide selection of drinks, live music, and a welcoming atmosphere with attentive service. The recovered profile distills these characteristics, focusing on the pub’s appeal for users who enjoy diverse beer options, live music, and a relaxed setting. This alignment suggests that our decoder reliably summarizes key attributes, albeit with a degree of abstraction.
Figure 6. A depiction of model prompt instruction.
Table 6. Experiments assessing decoder quality.
Data Amazon-books Yelp
BERTScore P R F1 P R F1
LightGCN 0.3266 0.2854 0.3064 0.3511 0.2238 0.2872
HICF 0.3498 0.3009 0.3257 0.3608 0.2398 0.3021
Ours 0.3665 0.3167 0.3418 0.3716 0.2453 0.3081

### v2

@@
-HARec: Hyperbolic Graph-LLM Alignment for Exploration and Exploitation in Recommender Systems
-Abstract.
-Modern recommendation systems often create information cocoons, limiting users’ exposure to diverse content.
-To enhance user experience, a crucial challenge is developing systems that can balance content exploration and exploitation, allowing users to adjust their recommendation preferences.
-Intuitively, this balance can be achieved through a tree-structured representation, where depth search facilitates exploitation and breadth search enables exploration.
-However, current works face two challenges to achieve this target: (1) Euclidean methods fail to fully capture hierarchical structures and lack flexibility in balancing exploration-exploitation, while (2) hyperbolic approaches, despite better hierarchical modeling, suffer from insufficient semantic alignment due to their reliance on Euclidean text encoders.
-To address these challenges, we propose HARec, a hyperbolic representation learning framework that jointly aligns user-item collaborative information with textual descriptions in hyperbolic space.
-Our framework introduces two key technique novelty: (1) a hierarchical-aware graph-llm alignment mechanism that enables better hierarchical representation, and (2) a hyperbolic hierarchical tree structure that facilitates user-adjustable exploration-exploitation trade-offs.
-Extensive experiments demonstrate that HARec consistently outperforms both Euclidean and hyperbolic baselines, achieving up to 5.49% improvement in utility metrics and 11.39% increase in diversity metrics.
-Recommender Systems, Representation Learning, Hyperbolic Space
-Authors affiliated with Snap Inc. served in advisory roles only for this work.
-Corresponding author
-copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; Aug 3–Aug 7, 2025; Woodstock, NYccs: Information systems Recommender systemsccs: Information systems Collaborative filteringccs: Mathematics of computing Geometric topologyccs: Information systems Personalization
-1. INTRODUCTION
-As online platforms grow exponentially, users are overwhelmed by an influx of information, resulting in information cocoons that restrict their exposure to diverse and novel content (Kunaver and Požrl, 2017; Castells et al., 2021).
+Breaking Information Cocoons: A Hyperbolic Graph-LLM Framework for Exploration and Exploitation in Recommender Systems
+Abstract
+Modern recommender systems often create information cocoons, restricting users’ exposure to diverse content.
+A key challenge lies in balancing content exploration and exploitation while allowing users to adjust their recommendation preferences.
+Intuitively, this balance can be modeled as a tree-structured representation, where depth search facilitates exploitation and breadth search enables exploration.
+However, existing approaches face two fundamental limitations: Euclidean methods struggle to capture hierarchical structures, while hyperbolic methods, despite their superior hierarchical modeling, lack semantic understanding of user and item profiles and fail to provide a principled mechanism for balancing exploration and exploitation.
+To address these challenges, we propose HERec, a hyperbolic graph-LLM framework that effectively balances exploration and exploitation in recommender systems.
+Our framework introduces two key innovations: (1) a hierarchical-aware graph-LLM mechanism that jointly aligns textual descriptions with user-item collaborative information in hyperbolic space, and (2) a hierarchical representation structure that enables user-adjustable exploration-exploitation trade-offs.
+Extensive experiments demonstrate that HERec consistently outperforms both Euclidean and hyperbolic baselines, achieving up to 5.49% improvement in utility metrics and 11.39% increase in diversity metrics, effectively mitigating information cocoons.
+1 INTRODUCTION
+As online platforms grow in size dramatically, users are overwhelmed by an influx of information, creating information cocoons that restrict their exposure to diverse and novel content (Kunaver & Požrl, 2017; Castells et al., 2021).
 Traditional recommendation approaches have primarily relied on collaborative filtering and interaction data (He et al., 2020; Koren et al., 2009), where users with similar historical preferences are expected to share future interests (Chen et al., 2021).
 However, these methods face two critical challenges in improving user experience.
-First, they often fail to capture the rich semantic relationships between items and evolving user preferences.
-For example, a jazz enthusiast might wish to explore rock music, yet current models struggle to understand these semantic connections and adapt to such evolving preferences (as illustrated in Figure 1).
-Second, existing approaches inadequately capture the underlying hierarchical structure inherent in user-item networks (Zheng et al., 2017; Unger and Tuzhilin, 2020), a critical factor for improving performance and personalization.
-The emergence of large language models (LLMs) offers a solution to the limitations of semantic understanding.
-However, integrating LLM-derived semantic features with traditional collaborative signals remains challenging, especially when capturing the hierarchical structure of user preferences and item relationships.
-Hyperbolic space has emerged as a promising solution to these issues.
-It is a powerful tool for modeling tree-like structures and power-law distributions (Krioukov et al., 2010; Sun et al., 2021; Yang et al., 2022b, a), where its exponential volume growth naturally aligns with hierarchical, scale-free structures (Ganea et al., 2018; Chami et al., 2019; Liu et al., 2019; Zhang et al., 2021; Yang et al., 2023).
+First, existing models struggle to capture the underlying hierarchical structure inherent in user-item networks (Zheng et al., 2017; Unger & Tuzhilin, 2020), a critical factor for enhancing performance and personalization.
+Second, these models fail to effectively balance the exploration-exploitation trade-off.
+For instance, a jazz enthusiast may want to explore rock music, but current models primarily focus on exploitation—recommending only familiar genres—without adequately supporting exploration (as illustrated in Figure 1).
+Hyperbolic space has emerged as a promising approach for modeling hierarchical structures and power-law distributions (Krioukov et al., 2010; Yang et al., 2022b; Sun et al., 2021; Yang et al., 2022a), where its exponential volume growth naturally aligns with hierarchical, scale-free structures (Ganea et al., 2018; Chami et al., 2019; Liu et al., 2019; Zhang et al., 2021; Yang et al., 2023).
 In hyperbolic space, distances increase exponentially from the origin, allowing popular items or exploratory users to be positioned closer to the origin, while niche items or users with focused interests spread toward the boundaries.
-This spatial configuration enables hyperbolic graph neural network (GNN) models to perceive hierarchical relationships effectively, enhancing recommendation performance (Sun et al., 2021; Yang et al., 2022a, b).
-However, current hyperbolic recommendation systems still face two key limitations.
-First, hyperbolic approaches often lack proper semantic alignment due to their dependence on Euclidean text encoders, which struggle to integrate within hyperbolic space.
-This mismatch leaves existing hyperbolic models vulnerable to noise in collaborative information, as they fail to capture semantic insights effectively, impacting their robustness and accuracy.
-Second, hyperbolic recommender systems lack a systematic approach to balancing exploration (introducing new content) and exploitation (focusing on existing preferences) (Wilson et al., 2021).
-This limitation hinders their adaptability to users’ evolving interests and reduces the diversity and relevance of recommended items.
+This spatial configuration enables hyperbolic graph neural network (GNN) models to perceive hierarchical relationships effectively, enhancing recommendation performance (Yang et al., 2022a; Sun et al., 2021; Yang et al., 2022b).
+Despite these advantages, current hyperbolic recommender systems still face key limitations.
+First, they struggle with a lack of semantic understanding of user and item profiles.
+While large language models (LLMs) offer a potential solution, their text encoders operate in Euclidean space, making it challenging to integrate LLM-derived semantic features with hyperbolic collaborative signals.
+This mismatch leaves existing hyperbolic models vulnerable to noise in collaborative information, as they fail to capture semantic insights effectively.
+Second, hyperbolic recommender systems lack a principled approach to balancing exploration (introducing new content) and exploitation (focusing on existing preferences) (Wilson et al., 2021).
+This limitation reduces their ability to adapt to the exploration-exploitation trade-off, ultimately diminishing both the utility and diversity of recommendations.
 Proposed Methods.
-To address these challenges, we propose HARec, a hyperbolic representation learning framework that facilitates exploration-exploitation through an intuitive tree-structured representation, where depth search enables exploitation and breadth search allows exploration.
-Our framework introduces two key components: (1) a hierarchical-aware graph-llm alignment mechanism that enhances semantic understanding and reduces noise sensitivity by jointly aligning textual descriptions with user-item collaborative information in hyperbolic space, and (2) a hyperbolic hierarchical tree structure that enables user-adjustable exploration-exploitation trade-offs.
-This tree structure organizes preferences hierarchically, with higher layers representing broader interests, allowing systematic navigation between exploration and exploitation.
-To validate the effectiveness of HARec, we conducted extensive experiments across multiple dimensions.
-Our model consistently outperformed baselines in both utility and diversity, marking a significant achievement as it is the first model to excel in both aspects simultaneously.
-To further highlight the benefits of our hyperbolic alignment framework, we compared HARec with feature-empowered baselines, where it demonstrated superior performance.
-Additionally, we conducted detailed analysis on head and tail items, showing that HARec effectively boosts recommendations for tail items, thus enhancing diversity.
-An ablation study, along with a decoder module analysis, confirms HARec’s capability to leverage both semantic meaning and collaborative information, providing deeper insights into the advantages of our approach.
+To address these challenges, we propose HERec, a hyperbolic graph-LLM framework designed to balance exploration and exploitation in recommender systems.
+Our approach introduces two key innovations: (1) a hierarchical-aware graph-LLM alignment mechanism, which jointly aligns textual descriptions with user-item collaborative information in hyperbolic space, supported by theoretical gradient analysis demonstrating its advantages, and (2) a novel hierarchical representation structure that enables user-adjustable exploration-exploitation trade-offs.
+Unlike conventional approaches that require manually defined hierarchies, we propose a hyperparameter-free clustering mechanism theoretically optimized by Dasgupta’s cost (Dasgupta, 2016).
+This approach automatically discovers hierarchical structures without requiring predefined hyperparameters, allowing for a more adaptive and principled organization of user preferences.
+To validate the effectiveness of HERec, we conducted extensive experiments across multiple dimensions.
+Our model consistently outperformed baselines in both utility and diversity metrics, marking a significant achievement as it is the first model to excel in both aspects simultaneously.
+For a fair comparison, we enhanced Euclidean baselines with feature information, yet our model still demonstrated superior performance, validating the advantage of our designed alignment in hyperbolic space.
+Additionally, we conducted a detailed analysis on head and tail items, showing that HERec effectively boosts recommendations for tail items, thus enhancing diversity.
+An ablation study, along with a hierarchical representation structure analysis and decoder module, confirms HERec’s capability to leverage both semantic meaning and collaborative information, providing deeper insights into the advantages of our approach.
 Contributions.
-To summarize, our primary contribution are in three aspects.
-The first is the proposed HARec, a hyperbolic representation learning framework that effectively balances exploration and exploitation by integrating semantic and collaborative information, supported by theoretical analysis demonstrating the advantages of this integration in hyperbolic space.
-In addition, we design a novel mechanism enabling personalized control over the exploration-exploitation trade-off through a hierarchical tree structure in hyperbolic space, allowing users to dynamically adjust their recommendation preferences.
-Moreover, through extensive experiments evaluating both accuracy and diversity metrics, we demonstrate that HARec achieves state-of-the-art performance across both criteria - achieving up to 5.49% improvement in utility metrics and 11.39% increase in diversity metrics, making it the first model to excel in both criteria simultaneously.
-Figure 1. Exploration and exploitation in hyperbolic space.
-2. Related Work
-In this section, we review commonly used graph-based collaborative filtering approaches, large language models (LLMs) for recommendation, and hyperbolic representation learning methods.
+Our primary contributions focus on three key aspects: capturing collaborative signals, integrating semantic information, and modeling the underlying hierarchical structure.
+The proposed graph-LLM alignment framework unifies collaborative and semantic information in hyperbolic space, while a hierarchical representation structure organizes user preferences, allowing for dynamic adjustment between exploration and exploitation.
+Moreover, through extensive experiments, we demonstrate that HERec achieves state-of-the-art performance across both utility and diversity metrics, making it the first model to excel in both criteria simultaneously.
+To facilitate reproducibility and further research, we open-source our model implementation at: https://github.com/Martin-qyma/HERec.
+Figure 1: Exploration and exploitation in hyperbolic space.
+2 Related Work
+In this section, we review graph-based collaborative filtering, large language models (LLMs) for recommendation, and hyperbolic representation learning methods.
 Graph Collaborative Filtering.
 Graph Collaborative Filtering has emerged as a dominant approach in recommender systems, with graph neural networks (GNNs) playing a central role.
-Unlike traditional collaborative filtering techniques, such as matrix factorization (Koren et al., 2009), GNNs capitalize on the relational structure inherent in user-item interactions, enabling a more nuanced understanding of complex dependencies.
-Prominent graph-based methods, such as NGCF (Wang et al., 2019), SGL (Wu et al., 2021), and LightGCN (He et al., 2020), have been widely adopted and have consistently achieved state-of-the-art performance, significantly boosting recommendation accuracy.
-Nevertheless, most of these models are constrained to Euclidean space, which may fall short in capturing the underlying power-law distribution commonly seen in user-item networks.
+Unlike traditional collaborative filtering techniques, such as matrix factorization (Koren et al., 2009), GNNs capitalize on the relational structure in user-item interactions, enabling a more nuanced understanding of complex dependencies.
+Prominent graph-based methods, such as NGCF (Wang et al., 2019), SGL (Wu et al., 2021), and LightGCN (He et al., 2020), have been widely adopted and have consistently achieved state-of-the-art performance.
+Nevertheless, these models are constrained to Euclidean space, which may fall short in capturing the underlying power-law distribution in user-item networks.
 Large Language Models for Recommendation.
-Large language models (LLMs) are increasingly used in recommender systems due to their strong capability in understanding textual information.
-Some approaches leverage LLMs directly for various downstream recommendation tasks (Bao et al., 2023; Li et al., 2023; Lyu et al., 2023; Zheng et al., 2024; Zhang et al., 2023), while others utilize LLMs as an auxiliary tool for collaborative filtering (Ren et al., 2024; Xi et al., 2024; Wei et al., 2024).
-However, both approaches often require substantial inference time and are computationally inefficient.
-In our HARec, we use language models solely to enhance data quality, avoiding the need for repeated inference during each recommendation by relying on hyperbolic embeddings for the final output.
+Large language models (LLMs) are increasingly used in recommender systems for their strong capability in understanding textual information.
+Some approaches leverage LLMs directly for downstream recommendation tasks (Bao et al., 2023; Li et al., 2023; Lyu et al., 2023; Zheng et al., 2024; Zhang et al., 2023), while others utilize LLMs as an auxiliary tool for collaborative filtering (Ren et al., 2024; Xi et al., 2024; Wei et al., 2024).
+However, these approaches require substantial inference time and are computationally inefficient.
+In our HERec, we use language models exclusively to improve data quality, eliminating the need for repeated inference during each recommendation by utilizing hyperbolic embeddings for the final output.
 Hyperbolic Representation Learning.
@@
 While some approaches incorporate knowledge graphs as supplementary information (Tai et al., 2021; Du et al., 2022), these models primarily focus on refining hierarchical structures rather than achieving a deeper semantic understanding of each item or user.
-3. PRELIMINARIES
-Embeddings learned from the above graph collaborative filtering are by default in the Euclidean space, which is unable to capture tree-like structures.
+3 PRELIMINARIES
+Embeddings learned from graph collaborative filtering are by default in the Euclidean space, which is unable to capture tree-like structures.
 To address the problem, we introduce how to derive these embeddings in a hyperbolic space.
@@
 An $n$-dimensional Lorentz manifold with negative curvature $-1/\kappa$ $(\kappa>0)$ is described as the Riemannian manifold $(\mathbb{H}^{n}_{\kappa},g_{\mathcal{L}})$.
-Here, $\mathbb{H}^{n}_{\kappa}=\{x\in\mathbb{R}^{n+1}:\langle x,x\rangle_{\mathcal{L}}=-\kappa,x_{0}>0\}$, $g_{\mathcal{L}}=\eta$ with $\eta=\mathbf{I}_{n}$ except $\eta_{0,0}=-1$, and $\langle\cdot,\cdot\rangle_{\mathcal{L}}$ represents the Lorentzian inner product, which is defined as by
-(1) $\langle x,y\rangle_{\mathcal{L}}:=-x_{0}y_{0}+\sum_{i=1}^{n}x_{i}y_{i}.$
+Here, $\mathbb{H}^{n}_{\kappa}=\{x\in\mathbb{R}^{n+1}:\langle x,x\rangle_{\mathcal{L}}=-\kappa,x_{0}>0\}$, $g_{\mathcal{L}}=\eta$ with $\eta=\mathbf{I}_{n}$ except $\eta_{0,0}=-1$, and $\langle\cdot,\cdot\rangle_{\mathcal{L}}$ represents the Lorentzian inner product, which is defined
+$\langle x,y\rangle_{\mathcal{L}}:=-x_{0}y_{0}+\sum_{i=1}^{n}x_{i}y_{i}$ (1)
 A tangent space is an $n$-dimensional vector space that approximates Euclidean space to hyperbolic space.
 Specifically, given a point $x\in\mathbb{H}_{\kappa}^{n}$, there exists a tangent space $\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}$ that approximates $\mathbb{H}_{\kappa}^{n}$:
-(2) $\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}:=\{v\in\mathbb{R}^{n+1}:\langle v,x\rangle_{\mathcal{L}}=0\}.$
-The tangent space is defined to be orthogonal to $x$ with respect to the Lorentzian inner product, therefore making the approximation accurate at the reference point.
+$\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}:=\{v\in\mathbb{R}^{n+1}:\langle v,x\rangle_{\mathcal{L}}=0\}$ (2)
 We also need to define transitions between $\mathbb{H}_{\kappa}^{n}$ and $\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}$, which give rise to the exponential and logarithmic maps.
@@
 The exponential map $\exp_{x}:\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}\to\mathbb{H}_{\kappa}^{n}$ is given by $\exp_{x}(v)=\gamma(1)$:
-(3) $\exp_{x}^{\kappa}(v)=\cosh\left(\frac{\|v\|_{\mathcal{L}}}{\sqrt{\kappa}}\right)x+\sqrt{\kappa}\sinh\left(\frac{\|v\|_{\mathcal{L}}}{\sqrt{\kappa}}\right)\frac{v}{\|v\|_{\mathcal{L}}},$
-where $\|v\|_{\mathcal{L}}=\sqrt{\langle v,v\rangle_{\mathcal{L}}}$ denotes the Lorentzian norm of $v$. The logarithmic map $\log_{x}$ is the inverse of the exponential map $\exp_{x}$:
-(4) $\log_{x}^{\kappa}(y)=d_{\mathcal{L}}^{\kappa}(x,y)\frac{y+\frac{1}{\kappa}\langle x,y\rangle_{\mathcal{L}}x}{\|y+\frac{1}{\kappa}\langle x,y\rangle_{\mathcal{L}}x\|},$
-where $d_{\mathcal{H}}^{\kappa}(x,y)$ represents the distance between $x$ and $y$ in $\mathbb{H}^{n}_{\kappa}$:
-(5) $d_{\mathcal{H}}^{\kappa}(x,y)=\sqrt{\kappa}\,\text{arcosh}\left(-\langle x,y\rangle_{\mathcal{L}}/\kappa\right).$
+$\exp_{x}^{\kappa}(v)=\cosh\left(\frac{\|v\|_{\mathcal{L}}}{\sqrt{\kappa}}\right)x+\sqrt{\kappa}\sinh\left(\frac{\|v\|_{\mathcal{L}}}{\sqrt{\kappa}}\right)\frac{v}{\|v\|_{\mathcal{L}}}$ (3)
+where $\|v\|_{\mathcal{L}}=\sqrt{\langle v,v\rangle_{\mathcal{L}}}$ denotes the Lorentzian norm of $v$.
+The logarithmic map $\log_{x}$ is the inverse of the exponential map $\exp_{x}$, and the distance between $x$ and $y$ in $\mathbb{H}^{n}_{\kappa}$ is:
+$\log_{x}^{\kappa}(y)=d_{\mathcal{L}}^{\kappa}(x,y)\frac{y+\frac{1}{\kappa}\langle x,y\rangle_{\mathcal{L}}x}{\|y+\frac{1}{\kappa}\langle x,y\rangle_{\mathcal{L}}x\|}$ (4)
+where $d_{\mathcal{H}}^{\kappa}(x,y)$ is the distance between $x$ and $y$ in $\mathbb{H}^{n}_{\kappa}$:
+$d_{\mathcal{H}}^{\kappa}(x,y)=\sqrt{\kappa}\,\text{arcosh}\left(-\langle x,y\rangle_{\mathcal{L}}/\kappa\right)$ (5)
 We choose $\mathbf{o}:=\{\sqrt{\kappa},0,\ldots,0\}\in\mathbb{H}^{n}_{\kappa}$ as the reference point for all the operations, making the transitions reproducible.
@@
 Therefore, in the following discussions, we would omit $\kappa$ for simplicity.
-4. METHODOLOGY
-In this section, we present a comprehensive overview of our proposed model, HARec.
+4 METHODOLOGY
+In this section, we present a comprehensive overview of our proposed model, HERec.
 Our approach not only aims to deliver accurate recommendations but also addresses users’ growing demand for more diverse suggestions.
@@
 The overall framework is in Figure 2.
-Figure 2.
-The overall architecture of HARec.
-(i) Hyperbolic Graph Collaborative Filtering: Encodes collaborative information using hyperbolic GNNs; (ii) Hyperbolic Alignment: Aligns semantic and collaborative information within hyperbolic space; (iii) Hyperbolic Hierarchical Structure: Builds hierarchy tree from hyperbolic embeddings.
-4.1. Hyperbolic Graph Collaborative Filtering
-Similar to Euclidean space, hyperbolic graph collaborative filtering captures high-order dependencies between users and items via a message aggregation mechanism.
+Figure 2: The overall architecture of HERec.
+(i) Hyperbolic Graph Collaborative Filtering: Encodes collaborative information using hyperbolic GNNs; (ii) Hyperbolic Alignment: Aligns semantic and collaborative information within hyperbolic space; (iii) Hierarchical Representation Structure: Builds hierarchy structure from hyperbolic embeddings.
+4.1 Hyperbolic Graph Collaborative Filtering
+Similar to Euclidean space, hyperbolic graph collaborative filtering captures user-item dependencies via message passing mechanism.
 In our method, we use the Lorentz representation for both users and items.
@@
 We adopt the Gaussian sampling method for the initialization of embeddings in Euclidean space, after that, by injecting them into the tangent space of the reference point, we derive the initial hyperbolic embeddings of both users and items:
-(6) $z_{u}^{0}=(0,e_{u}),\quad\qquad z_{i}^{0}$ $=(0,e_{i})$
+$z_{u}^{0}=(0,e_{u}),\quad\qquad z_{i}^{0}$ $=(0,e_{i})$ (6)
 $h_{u}^{0}=\exp_{\mathbf{o}}(z_{u}^{0}),\quad\quad h_{i}^{0}$ $=\exp_{\mathbf{o}}(z_{i}^{0})$
@@
 Given user $u$, item $i$, as well as their neighborhood $\mathcal{N}_{u},\mathcal{N}_{i}$, their embeddings at layer $l$ in the tangent space are updated as follows:
-(7) $z_{u}^{l}=z_{u}^{l-1}+\sum_{i\in\mathcal{N}_{u}}\frac{1}{|\mathcal{N}_{u}|}z_{i}^{l-1},\quad\quad z_{i}^{l}=z_{i}^{l-1}+\sum_{u\in\mathcal{N}_{i}}\frac{1}{|\mathcal{N}_{i}|}z_{u}^{l-1}$
+$z_{u}^{l}=z_{u}^{l-1}+\sum_{i\in\mathcal{N}_{u}}\frac{1}{|\mathcal{N}_{u}|}z_{i}^{l-1},\ z_{i}^{l}=z_{i}^{l-1}+\sum_{u\in\mathcal{N}_{i}}\frac{1}{|\mathcal{N}_{i}|}z_{u}^{l-1}$ (7)
 where $|\mathcal{N}_{u}|$ and $|\mathcal{N}_{i}|$ represent the number of neighbors of $u$ and $i$, respectively. The final tangent space embedding is derived by summing up embeddings at each layer:
-(8) $z_{u}=\sum_{l}z_{u}^{l},\quad\quad z_{i}=\sum_{l}z_{i}^{l}$
+$z_{u}=\sum_{l}z_{u}^{l},\quad\quad z_{i}=\sum_{l}z_{i}^{l}$ (8)
 Since $z_{u}$ and $z_{i}$ are on the tangent space, we still need to project it into hyperbolic space to derive the hyperbolic embeddings:
-(9) $h_{u}=\exp_{\mathbf{o}}(z_{u}),\quad\quad h_{i}=\exp_{\mathbf{o}}(z_{i})$
+$h_{u}=\exp_{\mathbf{o}}(z_{u}),\quad\quad h_{i}=\exp_{\mathbf{o}}(z_{i})$ (9)
 Hyperbolic Prediction Function.
 Through hyperbolic messaging passing, we’ve captured high-order relational information into hyperbolic embeddings.
-Hyperbolic distance is now used to define a prediction function, which is used to predict the possibility that a user would interact with an item:
-(10) $p(u,i)=\frac{1}{d_{\mathcal{H}}(h_{u},h_{i})}$
-Hyperbolic Margin Ranking Loss.
-Bayesian Personalized Ranking (BPR) loss has been widely applied in recommender systems (Rendle et al., 2012), aiming to pull together user-item pairs with recorded interactions while pushing apart those without.
-In our model, we adopt the hyperbolic margin ranking loss (Sun et al., 2021), designed to prevent the undesirable collapse seen in Euclidean space by adding an additional margin to the original BPR loss:
-(11) $\ell_{margin}(u,i,j)=\max(d_{\mathcal{H}}^{2}(h_{u},h_{i})-d_{\mathcal{H}}^{2}(h_{u},h_{j})+m,0),$
+Hyperbolic distance is now used to define a prediction function, which predicts the possibility that a user would interact with an item:
+$p(u,i)=\frac{1}{d_{\mathcal{H}}(h_{u},h_{i})}$ (10)
+Hyperbolic Margin Ranking and Informative Mechanism.
+To address the limitations of Euclidean-based approaches, we adopt the hyperbolic margin ranking loss (Sun et al., 2021), an extension of Bayesian Personalized Ranking (BPR) (Rendle et al., 2012).
+This loss pulls user-item pairs with observed interactions closer while pushing negative pairs apart, preventing the collapse seen in Euclidean space by incorporating a margin:
+$\ell_{m}(u,i,j)=\max(d_{\mathcal{H}}^{2}(h_{u},h_{i})-d_{\mathcal{H}}^{2}(h_{u},h_{j})+m,0)$ (11)
 where $d_{\mathcal{H}}$ represents hyperbolic distance, $m$ is the margin, $(u,i)$ is an observed interaction, and $(u,j)$ is a sampled negative pair.
-Hyperbolic Informative Mechanism.
-We further integrate a geometric-aware mechanism (Yang et al., 2022a) to enhance the pull-push process, consisting of two parts: Hyperbolic-aware Margin Learning (HAML) and Hyperbolic Informative Negative Sampling (HINS).
-Given that distances in hyperbolic space grow exponentially as embeddings deviate from the origin, we aim to assign a larger margin when user-item pairs are near the origin.
-HAML addresses this issue by dynamically adapting the margin as follows:
-(12) $m^{\mathcal{H}}_{ui}$ $=\mathrm{sigmoid}(\delta),$
+To further refine this pull-push process, we integrate a geometric-aware mechanism (Yang et al., 2022a) combining Hyperbolic-aware Margin Learning (HAML) and Hyperbolic Informative Negative Sampling (HINS).
+HAML dynamically adjusts the margin based on hyperbolic geometry, assigning a larger margin to pairs near the origin:
+$m^{\mathcal{H}}_{ui}$ $=\mathrm{sigmoid}(\delta)$ (12)
 $\delta$ $=\frac{d_{\mathcal{H}}^{2}(e_{u},\mathbf{o})+d_{\mathcal{H}}^{2}(e_{i},\mathbf{o})-d_{\mathcal{H}}^{2}(e_{u},e_{i})}{e_{u,0}\cdot e_{i,0}}$
-where $e_{u,0}$ and $e_{i,0}$ represent the zeroth coordinate elements of $e_{u}$ and $e_{i}$, respectively, corresponding to the embeddings norms in the hyperbolic space.
-HINS is an algorithm that samples negative nodes close to the positive item, yielding informative negative samples for both head and tail items.
-This is achieved by iteratively calculating the distances between the positive item and all candidate negative samples until the nearest one is identified.
-4.2. Semantic Representations Generation
-In this subsection, we introduce a paradigm for distilling useful information from raw text descriptions of items and users into dense semantic embeddings, which will be used later in our alignment framework that derives the final embeddings, integrating both collaborative information and semantic insights.
-Inspired by prior research (Ma et al., 2024; Ren et al., 2024) that leverages the power of large language models (LLMs) to extract useful user/item descriptions from raw, noisy textual data, we adopt a similar framework to generate user/item profiles with rich textual descriptions:
-(13) $\mathcal{P}_{i}=\textit{LLMs}(\mathcal{S}_{i},\mathcal{T}_{i})$
-where $\mathcal{S}_{i}$ represents system prompts to the LLMs, $\mathcal{T}_{i}$ denotes raw textual information, and $\mathcal{P}_{i}$ is the resulting user/item profile, consisting of sentence-level descriptions.
-While we’ve derived textual information for users/items, it cannot be directly fused with the embeddings generated from hyperbolic collaborative filtering.
-To integrate both types of knowledge, we encode the textual information into embeddings using pretrained text encoder models, such as BERT (Devlin, 2018):
-(14) $e_{i}=\textit{Encoder}(\mathcal{P}_{i})$
-4.3. Hyperbolic Alignment
-Given two sets of embeddings: hyperbolic embeddings with collaborative information and semantic embeddings with textual information, we aim to maximize the mutual information between them, creating unified embeddings that integrate both types of knowledge.
-To achieve this, we introduce a hyperbolic alignment loss, which mitigates the effects of noisy signals in representation learning and improves recommendations for tail items by aligning information within the same hyperbolic space.
-A theoretical proof of the advantages of alignment in hyperbolic space over Euclidean space is provided in the next subsection.
+where $e_{u,0}$ and $e_{i,0}$ are the zeroth coordinate elements of $e_{u}$ and $e_{i}$, respectively, corresponding to the embeddings norms in the hyperbolic space.
+HINS enhances negative sampling by prioritizing candidates close to the positive item, iteratively identifying the most informative negative samples.
+This method leverages hyperbolic geometry to effectively model head and tail items, improving utility performance in recommendations.
+4.2 Semantic Representations Generation
+In this subsection, we introduce a paradigm for distilling useful information from raw text descriptions of items and users into dense semantic embeddings, which will be used in our alignment framework to derive the final representation.
+Inspired by prior research (Ma et al., 2024; Ren et al., 2024) that leverages the power of large language models (LLMs) to extract useful user/item descriptions from raw, noisy textual data, we adopt a similar framework to generate user/item profiles with rich textual descriptions: $\mathcal{P}_{i}=\textit{LLMs}(\mathcal{S}_{i},\mathcal{T}_{i})$, where $\mathcal{S}_{i}$ represents system prompts to the LLMs, $\mathcal{T}_{i}$ denotes raw textual information, and $\mathcal{P}_{i}$ is the resulting user/item profile, consisting of sentence-level descriptions.
+While we’ve derived textual information for users/items, it cannot be directly fused with the embeddings generated from hyperbolic collaborative filtering due to discrepancies in embedding geometry.
+To integrate both types of knowledge, we encode the textual information in embeddings using pre-trained text encoder models: $e_{i}=\textit{Encoder}(\mathcal{P}_{i})$.
+While LLM embeddings could be an alternative, they often introduce higher computational overhead and may not be optimized for alignment with hyperbolic representations, making encoders a more practical choice.
+4.3 Hyperbolic Alignment
+We propose maximizing mutual information between hyperbolic embeddings (collaborative information) and semantic embeddings (textual information) to create unified embeddings that integrate both knowledge types.
+To achieve this, we introduce a hyperbolic alignment loss, which reduces noise in representations and improves tail item recommendations by aligning information in the same hyperbolic space.
+A theoretical advantage of hyperbolic alignment over Euclidean alignment is discussed in the next subsection.
 Before calculating the alignment loss, two steps are necessary: 1) the semantic embeddings must be adjusted to match the dimensionality of the collaborative embeddings, and 2) they should be transformed into a common representation space.
@@
 We then project the semantic embeddings into hyperbolic space so they can align with the hyperbolic collaborative embeddings:
-(15) $s_{i}$ $=\exp_{x}(e_{i})$
-$s^{\prime}_{i}$ $=\textit{MLP}(s_{i})$
-where $\exp_{x}$ is defined as in 3, $e_{i}$ represents the semantic embeddings from the encoder, $s_{i}$ is the dimension-aligned semantic embedding, and $s^{\prime}_{i}$ denotes the final projected semantic embedding in hyperbolic space.
+$s_{i}=\exp_{x}(e_{i}),\quad s^{\prime}_{i}=\textit{MLP}(s_{i})$ (13)
+where $\exp_{x}$ is defined as in Equation 3, $e_{i}$ represents the semantic embeddings from the encoder, $s_{i}$ is the dimension-aligned semantic embedding, and $s^{\prime}_{i}$ denotes the final projected semantic embedding in hyperbolic space.
 The hyperbolic alignment loss is then calculated as:
-(16) $\ell_{align}(i)=d_{\mathcal{H}}^{2}(h_{i},\textbf{s}^{\prime}_{i})$
+$\ell_{align}(i)=d_{\mathcal{H}}^{2}(h_{i},s^{\prime}_{i})$ (14)
 where $d_{\mathcal{H}}$ represents the hyperbolic distance, and $h_{i}$ is the hyperbolic embedding with collaborative information.
-4.4. Gradient Analysis of Semantic Alignment.
-To further explore the impact of aligning semantic embeddings within the hyperbolic model, we analyze the gradient behavior with respect to nodes of varying norms in Proposition 4.1.
+4.4 Gradient Analysis of Hyperbolic Alignment
+To further explore the impact of aligning semantic embeddings within the hyperbolic model, we analyze the gradient behavior with respect to nodes of varying norms in Proposition 4.1, proof can be found in Appendix A.
 This provides insight into the hierarchical preservation properties of the hyperbolic space, which are absent in Euclidean space.
 Proposition 4.1.
-The gradient magnitude of semantic alignment in hyperbolic space $\|\nabla_{\mathbf{x}}d_{\mathcal{H}}\|\approx\frac{\|\hat{\mathbf{x}}-\hat{\mathbf{y}}\|}{\|\mathbf{x}\|(1-\cos\theta)}$ enables adaptive updates for hierarchical preservation, while Euclidean space maintains constant magnitude $\|\nabla_{\mathbf{x}}d_{E}\|=1$, where the $d_{\mathcal{H}}$ and $d_{\mathcal{E}}$ represent the hyperbolic distance and Euclidean distance, respectively.
-Hyperbolic Space. Suppose $\mathbf{x}$ is a hyperbolic embedding derived from  4.1, and $\mathbf{y}$ is the semantic embedding from  4.2. The hyperbolic distance is given by:
-(17) $d_{\mathcal{H}}(\mathbf{x},\mathbf{y})=\operatorname{arccosh}(z)$
-where $z=-\langle\mathbf{x},\mathbf{y}\rangle_{\mathcal{L}}=x_{0}y_{0}-\sum_{i=1}^{n}x_{i}y_{i}$.
-For large $\|\mathbf{x}\|$ and $\|\mathbf{y}\|$, given $x_{0}=\sqrt{1+\|\mathbf{x}\|^{2}}$ and $y_{0}=\sqrt{1+\|\mathbf{y}\|^{2}}$, we adopt the following approximations for further calculation:
-(18) $x_{0}\approx\|\mathbf{x}\|,\quad y_{0}\approx\|\mathbf{y}\|,\quad\sqrt{z^{2}-1}\approx z$
-(19) $z=x_{0}y_{0}-\sum_{i=1}^{n}x_{i}y_{i}\approx\|\mathbf{x}\|\|\mathbf{y}\|(1-\cos\theta)$
-Let $x_{j}=\|\mathbf{x}\|\hat{x}_{j}$ and $y_{j}=\|\mathbf{y}\|\hat{y}_{j}$ (where $\hat{x}_{j}$ represents the direction of $x_{j}$). The partial derivative of $z$ with respect to $x_{j}$ is:
-(20) $\frac{\partial z}{\partial x_{j}}=\frac{x_{j}}{x_{0}}y_{0}-y_{j}\approx\frac{x_{j}}{\|\mathbf{x}\|}\|\mathbf{y}\|-y_{j}=\|\mathbf{y}\|(\hat{x}_{j}-\hat{y}_{j})$
-Thus, the gradient norm of $d_{\mathcal{H}}$ for $\mathbf{x}$ can be approximated by:
-(21) $\|\nabla_{\mathbf{x}}d_{\mathcal{H}}\|=\frac{\|\nabla_{\mathbf{x}}z\|}{\sqrt{z^{2}-1}}\approx\frac{\|\nabla_{\mathbf{x}}z\|}{z}=\frac{\|\mathbf{y}\|\|\hat{\mathbf{x}}-\hat{\mathbf{y}}\|}{\|\mathbf{x}\|\|\mathbf{y}\|(1-\cos\theta)}=\frac{\|\hat{\mathbf{x}}-\hat{\mathbf{y}}\|}{\|\mathbf{x}\|(1-\cos\theta)}$
-This result indicates that nodes with large norms undergo smaller gradient updates, preserving local structures, while nodes with smaller norms adjust more, refining global hierarchical relationships and capturing hierarchical structure effectively.
-Euclidean Space. For comparison, we examine the gradient behavior in Euclidean space. The Euclidean distance is defined as:
-(22) $f(\mathbf{x})=d_{E}(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|$
-The gradient of $f(\mathbf{x})$ with respect to $\mathbf{x}$ is:
-(23) $\nabla_{\mathbf{x}}f(\mathbf{x})=\frac{\mathbf{x}-\mathbf{y}}{\|\mathbf{x}-\mathbf{y}\|}=\frac{\mathbf{x}-\mathbf{y}}{d_{E}(\mathbf{x},\mathbf{y})}$
-which has a magnitude of:
-(24) $\|\nabla_{\mathbf{x}}f(\mathbf{x})\|=1$
-In summary, the analysis shows that hyperbolic space provides adaptive gradient magnitudes that depend on the norms of the nodes, effectively capturing and preserving hierarchical structures.
-In contrast, Euclidean space, with its constant gradient magnitude, cannot adapt to hierarchical nuances, making it less suitable for modeling hierarchical data in recommender systems.
-4.5. Hyperbolic Hierarchical Structure
-Figure 3. Hierarchical Structure Visualization.
-In this subsection, we introduce a mechanism that allows users to control the level of exploration in their recommendations by constructing a hierarchy tree based on the learned hyperbolic embeddings, an example is shown in Figure 3.
-These embeddings integrate both collaborative information and language semantics, enabling more nuanced recommendations.
+Let $\mathbf{x},\mathbf{y}$ denote embeddings in hyperbolic space $\mathcal{H}^{n}$ with corresponding unit direction vectors in space-like dimension $\hat{\mathbf{x}},\hat{\mathbf{y}}$, and let $\theta$ be the angle between them.
+The gradient magnitude of semantic alignment satisfies:
+$\|\nabla_{\mathbf{x}}d_{\mathcal{H}}(\mathbf{x},\mathbf{y})\|\approx\frac{\|\hat{\mathbf{x}}-\hat{\mathbf{y}}\|}{\|\mathbf{x}\|(1-\cos\theta)}.$ (15)
+This facilitates adaptive gradient updates that intrinsically preserve hierarchical structures.
+In contrast, Euclidean space exhibits constant gradient magnitude $\|\nabla_{\mathbf{x}}d_{\mathcal{E}}(\mathbf{x},\mathbf{y})\|=1$, where $d_{\mathcal{H}}$ and $d_{\mathcal{E}}$ denote hyperbolic and Euclidean distances respectively.
+This proposition highlights that hyperbolic space offers adaptive gradient magnitudes based on node norms.
+Nodes with large norms (fine-grained preferences, small positional changes cause significant distance changes) receive smaller gradient updates, ensuring precise distance adjustments and preserving local structures.
+Meanwhile, nodes with smaller norms (abstract preferences, positional changes have less impact on distances) adjust more, refining global relationships and effectively capturing the hierarchical structure.
+4.5 Hierarchical Representation Structure
+In this subsection, we design a mechanism that allows users to balance the exploration-exploitation trade-off by constructing a hierarchical representation structure based on the learned hyperbolic embeddings.
 Hierarchy Tree.
 The primary motivation for constructing a hierarchy tree is to uncover hidden structures in the data that are not explicitly reflected in the dataset.
-For instance, consider the Amazon-CD dataset: a CD might be categorized under genres like “folk music”, but it cannot be labeled as “folk music preferred by teenagers” or further refined to “folk music preferred by teenagers who live in large cities”.
+For instance, consider the Amazon-CD dataset: a CD might be categorized under genres like “folk music”, but it cannot be labeled as “folk music preferred by teenagers”.
 However, in reality, such nuanced preferences exist within the collaborative information.
-When recommending based solely on past interactions, we may overlook these hidden layers of structure that are implicit in the dataset (e.g., “music” → “folk music” → “folk music preferred by teenagers” → “folk music preferred by teenagers who live in large cities”).
-As previously discussed, in hyperbolic space, the norm of the embeddings reflects the level of exploration, where depth search facilitates exploitation and breadth search enables exploration.
-Building on this concept, we use the learned embeddings to form the bottom layer of the hierarchy tree.
-We then apply K-Means clustering in hyperbolic space to derive higher-level cluster nodes, which act as pseudo clustering nodes within the hierarchy, representing group preferences.
-For example, if an item is classified as “folk music preferred by teenagers who live in large cities”, its parent cluster may represent “folk music preferred by teenagers”, while its sibling clusters could include “folk music preferred by teenagers who live in small towns”.
-This hierarchical clustering process continues iteratively until reaching the root, thereby creating a multi-level tree structure that captures different levels of exploration.
+When recommending based solely on past interactions, we may overlook these hidden layers of structure that are implicit in the dataset (e.g., “music” → “folk music” → “folk music preferred by teenagers”).
+To address this, we introduce a hyperbolic hierarchical clustering method based on hyperbolic embeddings to build the hierarchy tree.
+Specifically, we iteratively group embeddings into clusters and use the centroid of each cluster as the data point to be clustered in the upper layer, referred to as a pseudo-cluster node.
+However, the number of cluster nodes in each layer is not predefined and must be determined.
+Inspired by Dasgupta’s cost (Dasgupta, 2016), which encourages pushing edge cuts as far down the tree as possible, the optimal tree is required to be binary.
+Therefore, we fix the proportion of cluster nodes between layers as $k=2$.
+The details of the algorithm are presented in Algorithm 1.
+Algorithm 1 Hyperbolic Hierarchical Clustering
+  Input: Hyperbolic embeddings $X$, layer proportion $k$
+  Set the maximum number of layers $L=\log_{k}(|X|)$
+  Initialize the points of bottom layer $D_{L}=X$
+  for $l=L$ to $1$ (bottom to top) do
+     Randomly initialize cluster centroids in the hyperbolic space $M_{l}=\{\mu_{l}\ |\ \mu_{l}\in\mathbb{H}\}$, $|M_{l}|=|D_{l}|/k$
+     repeat
+        for each point $d_{l,i}\in D_{l}$ do
+           Assign $d_{i}$ to the cluster with the nearest centroid:
+           $c_{l,i}=\arg\min_{\mu\in M_{l}}d_{\mathcal{H}}(d_{i},\mu)$
+        end for
+        for each cluster $c_{l,i}\in C_{l}$ do
+           Update the centroid $\mu_{l,i}$ as the hyperbolic mean $m_{\mathcal{H}}$ of all points assigned to cluster $c_{l,i}$:
+           $\mu_{l,i}=m_{\mathcal{H}}(\{d\ |\ d\in C_{l,i}\})$
+        end for
+     until Centroids $M_{l}$ do not change (convergence)
+     Initialize the points of the upper layer $D_{l-1}=M_{l}$
+  end for
 Exploration-Exploitation Balance.
-To allow users control over the balance of exploration and exploitation, we introduce two adjustable parameters: temperature $\tau$ and hierarchy level $l$.
+To allow users to balance the exploration-exploitation trade-off, we introduce two adjustable parameters: temperature $\tau$ and hierarchy layer $l$.
 In the standard recommendation approach, items are recommended based on similarity to the user’s embedding, which may limit exploration.
 For users with focused preferences, this method may consistently recommend items that align too closely with their past choices, offering little opportunity to explore different categories.
-To address this issue, we modify the recommendation process by replacing a portion of the recommended items with alternatives from other branches of the hierarchy tree.
+To address this issue, we modify the recommendation process by replacing a portion of the recommended items with alternatives from other branches in the hierarchy tree.
 This modification gives users control over the balance between exploration and exploitation.
-Temperature $\tau$ allows users to specify the proportion of original recommendations to replace, while level $l$ determines the extent of exploration within the hierarchy.
-For example, if a user selects $\tau=0.5$ and $l=2$, half of the originally recommended items will be retained, while the other half will be replaced by items sampled from nodes two levels up in the hierarchy tree, with selections made from the connected bottom-layer nodes.
-5. EXPERIMENTS
-Table 1.
-Comprehensive model comparison based on utility and diversity metrics, where higher values signify superior performance.
+Temperature $\tau$ allows users to specify the proportion of original recommendations to replace, while layer $l$ determines the starting point for exploration.
+For example, suppose a user selects $\tau=0.5$ and $l=5$.
+In that case, half of the original recommendations are retained, and the other half is replaced with items sampled from its ancestor cluster at layer $5$ of the hierarchy tree.
+Complexity Analysis.
+To demonstrate the scalability and efficiency of our approach, we analyze the complexity of constructing and applying the hierarchical representation structure.
+Let $N$ denote the total number of users and items, and let $k$ be the number of recommendations per user.
+The hierarchy tree is built using iterative $k$-means clustering in hyperbolic space, where cluster nodes are assigned only to real node embeddings for efficiency.
+First, computing and storing pairwise distances requires $O(N^{2}/2)$ operations.
+Each $k$-means iteration has complexity $O(i)$, and with $O(\log N)$ hierarchical levels, the total complexity of constructing the hierarchy tree is:$O\left(\frac{N^{2}}{2}+i\log N\right)$.
+For recommendation retrieval, finding the top-$k$ recommendations involves nearest neighbor search in hyperbolic space, which runs in $O(k\log N)$.
+Adjusting recommendations by retrieving ancestor clusters at layer $l$ requires only tree traversal, which runs in $O(k)$.
+Overall, our hierarchical framework ensures efficient scaling while maintaining flexibility in recommendation adjustments.
+This design enables HERec to dynamically balance exploration and exploitation while ensuring scalability for large-scale recommender systems.
+5 EXPERIMENTS
+Table 1: Comprehensive model comparison based on utility and diversity metrics, where higher values signify superior performance.
 The best-performing results for each metric are highlighted in bold, with the second-best results underlined.
@@
 Ours 0.1002 0.0772 0.1598 0.0992 0.3185 12.466 0.8598 0.3543 12.769 0.8728
-Table 2. Comparison with feature-enhanced Euclidean baselines, with the best-performing results highlighted in bold.
+Table 2: Comparison with feature-enhanced Euclidean baselines, with the best-performing results highlighted in bold.
 Data Amazon-books Yelp Google-reviews
@@
 Ours 0.1682 0.1069 0.5754 12.2543 0.8758 0.1224 0.0762 0.3961 12.3290 0.8773 0.1598 0.0992 0.3543 12.7690 0.8728
-5.1. Experimental Settings
-Datasets.
-To evaluate our proposed model, we utilize three widely-used public datasets: Amazon-books, which contains user purchase behaviors within the book category on Amazon; Yelp, which records customer ratings and reviews for restaurants on Yelp; and Google-reviews, which includes user reviews and business metadata from Google Maps.
-Detailed statistics for these datasets , including the distribution of head and tail items, are provided in Table 3.
-The high proportion of tail users (T80) across these datasets highlights the prevalence of long-tail distributions, supporting the assumption of a power-law distribution in real-world data.
-Evaluation Metrics.
-We assess our recommended items using both utility and diversity metrics, demonstrating that our model achieves precise recommendations for users’ preference while enhancing diversity in the recommendation list.
-For utility, we adopt two widely-used ranking metrics: Recall and NDCG, which measure model effectiveness.
-To evaluate diversity, we employ three metrics: Distance Diversity (Ziegler et al., 2005; Zhang and Hurley, 2008), Shannon Entropy (Mehta et al., 2012), and Expected Popularity Complement (Vargas and Castells, 2011).
-Distance Diversity (Div) quantifies the variation among recommended items, encouraging diversity within the recommendation:
-(25) $Div(u)=\frac{\sum_{i\in R_{u}}\sum_{j\in R_{u},j\neq i}d(i,j)}{\text{Number of}(i,j)\text{ pairs}}$
-where $R_{u}$ is the recommendation list for user $u$, and $d(i,j)$ represents the distance between items $i$ and $j$.
-Number of $(i,j)$ pairs is calculated as $\binom{|R_{u}|}{2}=\frac{|R_{u}|(|R_{u}|-1)}{2}$.
-For fair comparison, we compute distances within the same representation space across models.
-Shannon Entropy (H) quantifies the unpredictability within the recommendations, with higher entropy signifying greater diversity:
-(26) $H=-\sum_{i\in\text{set}(\sum_{u}R_{u})}p(i)\log p(i)$
-where $\sum_{u}R_{u}$ denotes all recommended items across users (including duplicates), and $p(i)=\text{count}(i)/\sum_{i\in\text{set}(\sum_{u}R_{u})}\text{count}(i)$ is the proportion of item $i$ among all recommended items.
-Expected Popularity Complement (EPC) assesses the extent of bias towards popular items by promoting niche recommendations:
-(27) $EPC=1-\frac{1}{|\text{set}(\sum R_{u})|}\sum_{i\in\text{set}(\sum_{u}R_{u})}\frac{\text{pop}(i)}{\max_{i}(\text{pop}(i))}$
-where $N_{u}$ is the set of items user $u$ has interacted with, and $\text{pop}(i)=\text{count}(i)/\sum_{i\in\text{set}(\sum_{u}N_{u})}\text{count}(i)$ denotes item $i$’s popularity.
-Baselines. We compare our model with state-of-the-art baselines in Euclidean and Hyperbolic spaces. Euclidean models include:
-  • LightGCN (He et al., 2020): A lightweight graph recommender model by removing redundant neural modules in graph convolution.
-  • TAG-CF (Ju et al., 2024): Enhances collaborative filtering by performing a single message-passing step only at inference time.
-  • LightGCL (Cai et al., 2023): Uses singular value decomposition (SVD) for robust graph augmentation without data augmentations.
-  • SimGCL (Yu et al., 2022): Replaces graph augmentations in contrastive learning by introducing random noise during initialization.
-  • HCCF (Xia et al., 2022): Integrates hypergraph and contrastive learning, capturing both local and global user dependencies.
-Hyperbolic GNN models include:
-  • HGCF (Sun et al., 2021): Combines hypergraph learning with collaborative filtering using hyperbolic margin ranking loss.
-  • HICF (Yang et al., 2022a): Enhances head and tail item performance with hyperbolic geometry, improving utility for long-tail items.
-Implementation Details.
-We set embedding dimensions to 50 and a training batch size of 1024.
-Embeddings in hyperbolic space are optimized using Riemannian SGD (RSGD) (Bécigneul and Ganea, 2018; Sun et al., 2021), while the MLP adapter is optimized with Adam (Kingma, 2014).
-Training incorporates early stopping technique based on Recall@10 with 10 patience steps.
-For the hierarchical structure, we use a $1:8$ ratio for parent and child nodes.
-We generate user and item profiles with GPT-3.5-turbo and encode them with text-embedding-ada-002 (Neelakantan et al., 2022) for Amazon-books and Yelp, and BERT (Devlin, 2018) for Google-reviews.
-The profile decoder backbone is Llama 3.1-8B.
-All experiments are conducted on NVIDIA A40/A100 GPUs with 40GB memory.
-Table 3. Statistics of the experimental data.
+Table 3: Statistics of the experimental data.
 Dataset #User #Item #Interactions Density
@@
 […]
-5.2. Performance Comparison
-Overall Comparison.
-The performance comparison across all models is presented in Table 1.
-Our HARec consistently outperforms the baselines in both utility and diversity metrics, achieving up to $5.49\%$ improvement in utility metrics and $11.39\%$ increase in diversity metrics.
-Notably, most baselines struggle to achieve strong performance in both utility and diversity simultaneously.
-For instance, while HICF performs second-best in utility metrics on the Amazon-books dataset, outperforming other baselines by a significant margin (excluding our model), it lacks the same advantage in diversity metrics, where several baselines achieve better results.
-This highlights the superiority of our model, which achieves state-of-the-art performance in both utility and diversity: a feat not previously attained by any baseline.
-Baselines with Feature Enhancement.
-To ensure a fair comparison, we further enhance Euclidean baselines by integrating feature information.
-Specifically, we align semantic embeddings with the original Euclidean embeddings following the structure in Section 4.2, but in Euclidean space.
-The results are presented in Table 2, with enhanced baselines denoted by “+”.
-The data shows that our HARec still surpasses all baselines in both utility and diversity.
-Interestingly, while the feature-enhanced versions of the Euclidean baselines generally improve in utility metrics, they do not exhibit a stable improvement in diversity metrics.
-This aligns with our theoretical analysis in Section 4.5, which demonstrates the advantages of our alignment model in the hyperbolic space.
-Table 4. Layer-wise norms within the hierarchy structure.
-Data level-0 level-1 level-2 level-3 level-4
-Amazon-books 0.2388 0.4859 0.5143 0.5889 0.6773
-Yelp 0.1758 0.3500 0.4719 0.5590 0.6307
-Google-reviews 0.4374 0.8585 1.0682 1.1731 1.2479
-Figure 4. Ablation study on model variants.
-Table 5.
-Performance on H20 (head) and T80 (tail) items, with the top hyperbolic results highlighted in bold.
+Table 4: Performance on H20 (head) and T80 (tail) items, with the top hyperbolic results highlighted in bold.
 $\Delta_{\mathcal{H}}(\%)$ represents the improvement over the second-best hyperbolic baseline, which is underlined.
-Data Amazon-books Yelp Google-reviews
+Datasets Amazon-books Yelp Google-reviews
 Metrics Recall@20 NDCG@20 Recall@20 NDCG@20 Recall@20 NDCG@20
@@
 $\Delta_{\mathcal{H}}(\%)$ +3.59 +8.58 +2.38 +9.71 +1.89 +12.11 +0.36 +13.33 +0.70 +14.18 +1.21 +14.42
-Figure 5. Performance across different levels of exploration.
-5.3. Ablation Study
+5.1 Experimental Settings
+Datasets.
+To evaluate our proposed model, we utilize three widely-used public datasets: Amazon-books, which contains user purchase behaviors within the book category on Amazon; Yelp, which records customer ratings and reviews for restaurants on Yelp; and Google-reviews, which includes user reviews and business metadata from Google Maps.
+Detailed statistics, including the distribution of head and tail items, are provided in Table 3.
+The high proportion of tail users (T80) highlights the prevalence of long-tail distributions, supporting the assumption of a power-law distribution in real-world data.
+Evaluation Metrics.
+We assess our recommended items using both utility and diversity metrics, demonstrating that our model achieves precise recommendations for users’ preference while enhancing diversity in the recommendation.
+For utility, we adopt two widely used ranking metrics: Recall and NDCG, which measure the effectiveness of the model.
+To evaluate diversity, we employ three metrics: Distance Diversity (Div) (Ziegler et al., 2005; Zhang & Hurley, 2008), Shannon Entropy (H) (Mehta et al., 2012), and Expected Popularity Complement (EPC) (Vargas & Castells, 2011).
+Distance Diversity (Div) quantifies the variation among recommended items, encouraging diversity within the recommendation:
+$Div(u)=\frac{\sum_{i\in R_{u}}\sum_{j\in R_{u},j\neq i}d(i,j)}{\text{Number of}(i,j)\text{ pairs}}$ (16)
+where $R_{u}$ is the recommendation list for user $u$, and $d(i,j)$ represents the distance between items $i$ and $j$.
+Number of $(i,j)$ pairs is calculated as $\binom{|R_{u}|}{2}=\frac{|R_{u}|(|R_{u}|-1)}{2}$.
+For fair comparison, we compute distances within the same representation space across models.
+Shannon Entropy (H) quantifies the unpredictability within the recommendations, with higher entropy signifying greater diversity:
+$H=-\sum_{i\in\text{set}(\sum_{u}R_{u})}p(i)\log p(i)$ (17)
+where $\sum_{u}R_{u}$ denotes all recommended items across users (including duplicates), and $p(i)=\text{count}(i)/\sum_{i\in\text{set}(\sum_{u}R_{u})}\text{count}(i)$ is the proportion of item $i$ among all recommended items.
+Expected Popularity Complement (EPC) assesses the extent of bias towards popular items by promoting niche recommendations:
+$EPC=1-\frac{1}{|\text{set}(\sum R_{u})|}\sum_{i\in\text{set}(\sum_{u}R_{u})}\frac{\text{pop}(i)}{\max_{i}(\text{pop}(i))}$ (18)
+where $N_{u}$ is the set of items user $u$ has interacted with, and $\text{pop}(i)=\text{count}(i)/\sum_{i\in\text{set}(\sum_{u}N_{u})}\text{count}(i)$ denotes item $i$’s popularity.
+Baselines. We compare our model with state-of-the-art baselines in both Euclidean and Hyperbolic spaces:
+  • LightGCN (He et al., 2020): A lightweight graph convolution neural network model for recommendation by removing redundant neural modules in graph convolution.
+  • TAG-CF (Ju et al., 2024): A test-time augmentation framework that enhances recommendation by performing a single message-passing step only at inference time.
+  • LightGCL (Cai et al., 2023): A contrastive learning model that uses singular value decomposition (SVD) for robust graph augmentation without data augmentation.
+  • SimGCL (Yu et al., 2022): A contrastive learning model that replaces graph augmentations in contrastive learning by introducing random noise during initialization.
+  • HCCF (Xia et al., 2022): A collaborative filtering framework that integrates hypergraph and contrastive learning, capturing both local and global user dependencies.
+  • HGCF (Sun et al., 2021): A hyperbolic GNN model that combines hypergraph learning with collaborative filtering using hyperbolic margin ranking loss.
+  • HICF (Yang et al., 2022a): A hyperbolic recommendation model that enhances head and tail item performance with hyperbolic geometry, improving utility for long-tail items.
+Training Details.
+We set embedding dimensions to 50 and a training batch size of 1024.
+Embeddings in hyperbolic space are optimized using Riemannian SGD (RSGD) (Bécigneul & Ganea, 2018; Sun et al., 2021), while the MLP adapter is optimized with Adam (Kingma, 2014).
+Training incorporates early stopping technique based on Recall@10 with 10 patience steps.
+We generate user and item profiles using GPT-3.5-turbo and encode them with text-embedding-ada-002 (Neelakantan et al., 2022) for Amazon-books and Yelp, and BERT (Devlin, 2018) for Google-reviews.
+All experiments are conducted on NVIDIA A40/A100 GPUs with 40GB memory.
+Figure 3: Ablation study on model variants.
+5.2 Performance Comparison
+Overall Comparison.
+The performance comparison across all models is presented in Table 1.
+Our HERec consistently outperforms the baselines in both utility and diversity metrics, achieving up to $5.49\%$ improvement in utility metrics and $11.39\%$ increase in diversity metrics.
+Notably, most baselines struggle to achieve strong performance in both utility and diversity simultaneously.
+For instance, while HICF performs second-best in utility metrics on the Amazon-books dataset, it lacks the same advantage in diversity metrics, where several baselines achieve better results.
+This highlights the superiority of our model, which achieves state-of-the-art performance in both utility and diversity: a feat not previously attained by any baseline.
+Baselines with Feature Enhancement.
+To ensure a fair comparison, we further enhance Euclidean baselines by integrating feature information.
+Specifically, we align semantic embeddings with the original Euclidean embeddings following the structure in Section 4.2, but in Euclidean space.
+The results are presented in Table 2, with enhanced baselines denoted by “+”.
+The data shows that our HERec still surpasses all baselines in both utility and diversity.
+Interestingly, while the feature-enhanced versions of the Euclidean baselines generally improve in utility metrics, they do not exhibit a stable improvement in diversity metrics.
+This aligns with our theoretical analysis in Section 4.4, which demonstrates the advantages of our alignment in the hyperbolic space.
+Additionally, to validate the effectiveness of our alignment framework in capturing semantics, we designed a profile decoder module to reconstruct user and item profiles from the learned hyperbolic embeddings.
+The results demonstrate that profiles reconstructed from our model’s embeddings exhibit higher alignment with the originals, confirming the model’s ability to effectively capture underlying semantic meaning.
+Detailed methodology and experimental results are provided in Appendix B.
+Figure 4: Performance analysis across hierarchical structure layers.
+Figure 5: Layer-wise norms within the hierarchy structure.
+Performance on Head and Tail Items.
+To underscore the effectiveness of semantic alignment, particularly for cold-start (tail) items, we conducted experiments measuring utility performance on both head and tail items.
+In our context, head items (H20) refer to the top $20\%$ of items with the highest number of interactions in the training dataset, representing the most popular items.
+The remaining items are categorized as tail items (T80), which have limited interaction history, representing less frequently engaged items.
+For simplicity, we compare our model with all hyperbolic baselines and a representative Euclidean baseline (LightGCN).
+The results, displayed in Table 4, reveal significant improvements, especially for tail items.
+Compared with hyperbolic baselines without semantic alignment, our model shows a notably larger performance boost on tail items than on head items.
+This finding indicates that semantic alignment particularly enhances the recommendation of cold-start items, effectively mitigating the cold-start problem for items with limited or no prior interactions.
+5.3 Ablation Study
 We perform an ablation study to examine the contributions of two key components in our model: hyperbolic margin ranking loss and semantic alignment loss.
@@
 Figure 3 shows that our full model outperforms all ablation variants, highlighting its overall effectiveness.
-Models lacking both margin ranking and semantic alignment perform the worst, while removing either module results in decreased utility, emphasizing the importance of each component and their effective integration.
-5.4. Hierarchical Structure Analysis
-To validate the effectiveness of our hierarchical structure, which allows users to control their level of exploration within recommendations, we conducted a series of analyses.
-Our goal is to demonstrate how the hierarchical model provides a structured balance between exploration and exploitation, allowing users to adjust the generality of recommendations by themselves.
-First, we analyze the norms of node embeddings across hierarchical layers.
-As shown in Table 4, nodes in higher layers have consistently smaller norms.
-This observation supports our design intention: higher-layer nodes represent broader, generalized group preferences, aligning with findings in Section 4.5.
-This structure enables more abstract grouping as the layers ascend, thereby capturing progressively broader user interests.
-To further examine the hierarchy’s practical utility, we designed experiments that simulate varying user preferences in daily interactions.
-Using a temperature parameter $\tau=0.2$ to control recommendation randomness, we tested recommendation performance across different hierarchical levels, from layer 00 (original recommendation level) up to layer $3$.
-The results, visualized in Figure 5, reveal a clear pattern: as the level $l$ increases, utility metrics tend to decrease, while diversity metrics increase.
-This progression highlights the intended trade-off our model offers: recommendations from lower layers (e.g., layer 00) focus on utility and specific user interests, ideal for exploitation.
-In contrast, higher layers (e.g., layer $3$) prioritize diversity by presenting broader, less tailored options, thus encouraging exploration.
-This systematic change across levels provides users with a flexible mechanism to adjust recommendation focus: from niche, highly relevant items to a more varied, exploratory set—based on their current preference for exploration or exploitation.
-To underscore the effectiveness of semantic alignment, particularly for cold-start (tail) items, we conducted experiments measuring utility performance on both head and tail items.
-In our context, head items (H20) refer to the top $20\%$ of items with the highest number of interactions in the training dataset, representing the most popular items.
-The remaining items are categorized as tail items (T80), which have limited interaction history, representing less frequently engaged (cold-start) items.
-For simplicity, we compare our model with all hyperbolic baselines and a representative Euclidean baseline (LightGCN).
-The results, displayed in Table 5, reveal significant improvements, especially for tail items.
-Compared with hyperbolic baselines without semantic alignment, our model shows a notably larger performance boost on tail items than on head items.
-This finding indicates that semantic alignment particularly enhances the recommendation of cold-start items, effectively mitigating the cold-start problem for items with limited or no prior interactions.
-6. CONCLUSION
-In this work, we introduce HARec, a novel hyperbolic graph-text alignment framework that effectively balances exploration and exploitation by integrating semantic and collaborative information, with theoretical analysis supporting the advantages of this approach within hyperbolic space.
-Additionally, we propose a unique mechanism that enables personalized control over the exploration-exploitation trade-off through a hierarchical tree structure, allowing users to adjust recommendation preferences dynamically.
-Extensive experiments demonstrate that HARec achieves state-of-the-art performance, with up to a $5.49\%$ improvement in utility metrics and an $11.39\%$ increase in diversity metrics, making it the first model to excel in both criteria simultaneously, a significant achievement that has yet to be attained by any previous models.
-Appendix A Appendix
-In the supplementary materials, we provide additional details on our datasets and present several case studies illustrating the performance of our decoder module.
-A.1. Profile Decoder
-To further evaluate the expressiveness of our learned hyperbolic embeddings, we introduce a profile decoder designed to reconstruct the original profile from these embeddings, thereby broadening their application beyond recommendation tasks.
+Models lacking both margin ranking and semantic alignment perform the worst, while removing either module results in decreased utility, emphasizing the effectiveness of integration.
+5.4 Hierarchical Representation Structure Analysis
+We evaluated the effectiveness of our hierarchical representation structure through experiments simulating varying user preferences.
+Using a temperature parameter $\tau=0.5$ to control recommendation randomness, we analyzed performance across various hierarchical layers.
+As shown in Figure 4, utility metrics increase while diversity metrics decrease with larger layers.
+This trade-off aligns with the model’s design: larger layers focus on utility and specific interests for exploitation, while smaller layers prioritize diversity for exploration.
+This structure provides users with a flexible mechanism to adjust recommendations from niche, relevant items to broader, exploratory sets based on their preferences.
+A case study is provided in Appendix C.
+Additionally, to validate the hierarchical structure’s capability to capture varying degrees of group preferences, we analyzed the average norm of node embeddings across layers (Figure 5).
+The results reveal a consistent increase at larger layers (farther from the center in hyperbolic space).
+As previously mentioned, hyperbolic geometry ensures that popular items or exploratory users are positioned closer to the origin, while niche items or users are positioned toward the boundaries.
+This finding aligns seamlessly with the properties of hyperbolic space, validating the intended design of the hierarchical structure: higher-level (smaller-layer) nodes effectively encapsulate abstract and aggregated preferences, whereas lower-level (larger-layer) nodes focus on fine-grained, personalized preferences.
+6 CONCLUSION
+In this work, we present HERec, a hyperbolic graph-LLM framework that effectively balances exploration and exploitation in recommender systems.
+Our approach integrates semantic and collaborative information within hyperbolic space, leveraging the unique properties of hyperbolic geometry.
+Furthermore, we introduce a personalized mechanism for balancing the exploration-exploitation trade-off using a hierarchical representation structure, empowering users to adjust recommendation preferences flexibly.
+Extensive experiments demonstrate that our model achieves state-of-the-art performance, achieving up to a $5.49\%$ improvement in utility metrics and an $11.39\%$ increase in diversity metrics.
+Remarkably, HERec is the first framework to excel in both utility and diversity simultaneously, marking a significant advancement in the field.
+ACKNOWLEGDEMENTS
+This work was partially supported by the National Science Foundation (NSF) IIS Div Of Information & Intelligent Systems 2403317.
+We also gratefully acknowledge partial support from an Amazon Research Award, funding generously provided by Snap Research, and resources made available by the Yale Office of the Provost.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Appendix A Proof
+In this section, we provide a rigorous mathematical proof of Proposition 4.1, demonstrating the unique advantage of the aligning mechanism in the hyperbolic space.
+A.1 Gradient Analysis in Hyperbolic Space.
+Suppose $\mathbf{x}$ is a hyperbolic embedding derived from  4.1, and $\mathbf{y}$ is the corresponding semantic embedding from  4.2, the hyperbolic distance is given by:
+$d_{\mathcal{H}}(\mathbf{x},\mathbf{y})=\operatorname{arccosh}(z)$ (19)
+where $z=-\langle\mathbf{x},\mathbf{y}\rangle_{\mathcal{L}}=x_{0}y_{0}-\sum_{i=1}^{n}x_{i}y_{i}$.
+For large $\|\mathbf{x}\|$ and $\|\mathbf{y}\|$, given $x_{0}=\sqrt{1+\|\mathbf{x}\|^{2}}$ and $y_{0}=\sqrt{1+\|\mathbf{y}\|^{2}}$, we adopt the following approximations for further calculation:
+$x_{0}\approx\|\mathbf{x}\|,\quad y_{0}\approx\|\mathbf{y}\|,\quad\sqrt{z^{2}-1}\approx z$ (20)
+Thus, the Lorentz product can be approximated as:
+$z=x_{0}y_{0}-\sum_{i=1}^{n}x_{i}y_{i}\approx\|\mathbf{x}\|\|\mathbf{y}\|(1-\cos\theta)$ (21)
+Let $x_{j}=\|\mathbf{x}\|\hat{x}_{j}$ and $y_{j}=\|\mathbf{y}\|\hat{y}_{j}$ (where $\hat{x}_{j}$ represents the direction of $x_{j}$).
+The partial derivative of $z$ with respect to $x_{j}$ is calculated as follows:
+$\frac{\partial z}{\partial x_{j}}$ $=\frac{x_{j}}{x_{0}}y_{0}-y_{j}$ (22)
+$\approx\frac{x_{j}}{\|\mathbf{x}\|}\|\mathbf{y}\|-y_{j}$
+$=\|\mathbf{y}\|(\hat{x}_{j}-\hat{y}_{j})$
+Therefore, the gradient norm of $d_{\mathcal{H}}$ for $\mathbf{x}$ can be approximated by:
+$\|\nabla_{\mathbf{x}}d_{\mathcal{H}}\|$ $=\frac{\|\nabla_{\mathbf{x}}z\|}{\sqrt{z^{2}-1}}$ (23)
+$\approx\frac{\|\nabla_{\mathbf{x}}z\|}{z}$
+$=\frac{\|\mathbf{y}\|\|\hat{\mathbf{x}}-\hat{\mathbf{y}}\|}{\|\mathbf{x}\|\|\mathbf{y}\|(1-\cos\theta)}$
+$=\frac{\|\hat{\mathbf{x}}-\hat{\mathbf{y}}\|}{\|\mathbf{x}\|(1-\cos\theta)}$
+This result suggests that nodes with large norms experience smaller gradient updates, preserving local structures, while nodes with smaller norms undergo larger updates, refining global hierarchical relationships and effectively capturing hierarchical structures.
+This aligns with the properties of hyperbolic geometry, where small positional changes near the origin result in significant distance changes, and vice versa.
+Our dynamic gradient update mechanism is well-suited to this characteristic, enabling precise adjustments and enhancing the accuracy of the updates.
+A.2 Gradient Analysis Euclidean Space.
+For comparison, we examine the gradient behavior in Euclidean space. The Euclidean distance is defined as:
+$f(\mathbf{x})=d_{E}(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|$ (24)
+The gradient of $f(\mathbf{x})$ with respect to $\mathbf{x}$ is:
+$\nabla_{\mathbf{x}}f(\mathbf{x})=\frac{\mathbf{x}-\mathbf{y}}{\|\mathbf{x}-\mathbf{y}\|}=\frac{\mathbf{x}-\mathbf{y}}{d_{E}(\mathbf{x},\mathbf{y})}$ (25)
+which has a magnitude of:
+$\|\nabla_{\mathbf{x}}f(\mathbf{x})\|=1$ (26)
+The constant gradient magnitude demonstrates Euclidean space’s uniform update behavior, lacking intrinsic adaptation to hierarchical structures.
+This fundamental difference in gradient dynamics explains hyperbolic space’s superiority for hierarchical representation learning.
+Appendix B Profile Decoder
+Figure 6: A depiction of model prompt instruction.
+To further evaluate the effectiveness of hyperbolic alignment and the expressiveness of our learned hyperbolic embeddings, we introduce a profile decoder designed to reconstruct the original profile from these embeddings, thereby broadening their application beyond recommendation tasks.
 Our method begins by mapping the hyperbolic embeddings back into Euclidean space to facilitate alignment with traditional network layers.
@@
 To enhance training efficiency and minimize computational cost, we freeze the language model’s parameters, making the MoE module the only trainable component in the architecture.
-Details of the input prompt structure and processing flow are depicted in Figure 6.
-To facilitate seamless embedding integration, we introduce two additional tokens, ¡USER_EMBED¿ and ¡ITEM_EMBED¿, to the language model’s tokenizer.
-This addition enables these embeddings to be treated as single, coherent tokens, ensuring accurate tokenization and embedding interpretation.
-After passing through the positional layer, the placeholders are dynamically replaced by the adapted embeddings generated by the MoE, allowing the embeddings to retain context-specific information.
-A.2. Decoder Quality
+[…]
+B.1 Decoder Quality
 To assess the effectiveness of our decoder mechanism, we compare the embeddings from our model against those from Euclidean embeddings (LightGCN) and Hyperbolic embeddings (HICF) as inputs to the decoder.
 The results are shown in Table 5.
-We use BERTScore (Zhang et al., [n.
-d.]) to evaluate the semantic alignment of generated sentences with the original profile.
-Results indicate that our HARec outperforms both Euclidean and hyperbolic baselines, demonstrating that our model’s embeddings support recommendation and retain user/item profile information.
+We use BERTScore (Zhang et al., 2019) to evaluate the semantic alignment of generated sentences with the original profile.
+Results indicate that our HERec outperforms both Euclidean and hyperbolic baselines, demonstrating that our model’s embeddings support recommendation and retain user/item profile information.
 This finding underscores that our model captures collaborative information alongside recoverable language semantics: a capability not achieved by previous models.
-A.2.1. Decoder Case Study
+B.2 Decoder Case Study
 To vividly illustrate the semantic capacity of our embeddings, we present two case studies demonstrating the model’s ability to recover user and item profiles.
-————————————————————————————————————— Original user profile:
+ Original user profile:
 Based on the user’s interactions and reviews, this user is likely to enjoy unique dining experiences with a variety of options, trendy atmosphere with good drink and food selections, premium quality frozen yogurt with a wide variety of toppings, creative burgers and hot dogs, delicious Thai cuisine, and authentic and spicy Chinese cuisine served family-style.
@@
 They also value quality and freshness in their food.
-————————————————————————————————————— In this example, the profile recovered by our decoder accurately captures the user’s preferences for Thai and Chinese cuisines, broadly summarized as an affinity for Asian fusion flavours.
+In this example, the profile recovered by our decoder accurately captures the user’s preferences for Thai and Chinese cuisines, broadly summarized as an affinity for Asian fusion flavours.
 Additionally, our decoder identifies the user’s interest in diverse culinary experiences, indicating a propensity toward exploration.
-————————————————————————————————————— Original item profile:
+ Original item profile:
 Ceol Irish Pub in Reno is a popular bar and pub known for its happy hour deals, great service, and friendly atmosphere.
@@
 Users who enjoy a variety of beer options, live music, and a relaxed atmosphere would enjoy Ceol Irish Pub.
-————————————————————————————————————— The alignment between the original and recovered item profiles shows that our decoder effectively captures key thematic elements of Ceol Irish Pub.
+The alignment between the original and recovered item profiles shows that our decoder effectively captures key thematic elements of Ceol Irish Pub.
 The original profile emphasizes aspects such as happy hour, a wide selection of drinks, live music, and a welcoming atmosphere with attentive service.
@@
 This alignment suggests that our decoder reliably summarizes key attributes, albeit with a degree of abstraction.
-Figure 6. A depiction of model prompt instruction.
-Table 6. Experiments assessing decoder quality.
+Table 5: Experiments assessing decoder quality.
 Data Amazon-books Yelp
 BERTScore P R F1 P R F1
-LightGCN 0.3266 0.2854 0.3064 0.3511 0.2238 0.2872
-HICF 0.3498 0.3009 0.3257 0.3608 0.2398 0.3021
+Euclidean 0.3266 0.2854 0.3064 0.3511 0.2238 0.2872
+Hyperbolic 0.3498 0.3009 0.3257 0.3608 0.2398 0.3021
 Ours 0.3665 0.3167 0.3418 0.3716 0.2453 0.3081
+Appendix C Hierarchical Structure Case Study
+In this section, we present a case study examining the exploration-exploitation trade-off in the hyperbolic hierarchical structure.
+We randomly select a user and identify items that share different layers of the lowest common ancestor (LCA) with the user in the hierarchy tree.
+Ideally, a larger LCA layer indicates that the user and item were grouped earlier in the hierarchy, suggesting a stronger likelihood of interaction.
+Conversely, a smaller LCA layer implies that the user and item were grouped only after multiple iterations, indicating a weaker likelihood of interaction.
+ User Profile:
+The user enjoys books with supernatural elements, horror and paranormal romance.
+They also appreciate fast-paced action and suspenseful thrillers with political intrigue and espionage.
+The user likes complex characters dealing with dark themes.
+Item 1 Profile (LCA=10):
+The book ’Spirit Fighter (Son of Angels, Jonah Stone)’ would likely appeal to young readers who enjoy stories about supernatural beings and the fight between good and evil.
+It is a book that offers entertainment and inspiration, as well as a glimpse into the supernatural world.
+Item 2 Profile (LCA=5):
+Fans of contemporary romance novels will enjoy Barefoot Summer (A Chapel Springs Romance).
+This book combines heartwarming romance, relatable heartaches, and lighthearted humor to create a delightful and engaging story.
+It is ideal for those looking for an easy-to-read and entertaining summer novel.
+As demonstrated in the case study, the user shows a preference for supernatural elements, which aligns closely with the description of item 1, indicating a high likelihood of interaction.
+In contrast, while the user profile mentions a preference for dark themes, item 2 focuses on romance and heartwarming narratives, suggesting a low probability of user interest in item 2.
+This analysis aligns seamlessly with the intended design of our hierarchical tree, further validating the effectiveness of our model in capturing and leveraging nuanced user-item relationships.

### v3

@@
 To address these challenges, we propose HERec, a hyperbolic graph-LLM framework that effectively balances exploration and exploitation in recommender systems.
-Our framework introduces two key innovations: (1) a hierarchical-aware graph-LLM mechanism that jointly aligns textual descriptions with user-item collaborative information in hyperbolic space, and (2) a hierarchical representation structure that enables user-adjustable exploration-exploitation trade-offs.
+Our framework introduces two key innovations: (1) a semantic-enhanced hierarchical mechanism that aligns rich textual descriptions processed by large language models (LLMs) with collaborative information directly in hyperbolic space, allowing for more nuanced updates that respect the underlying hierarchical structure in user-item profiles; (2) an automatic hierarchical representation by optimizing Dasgupta’s cost, which discovers hierarchical structures without requiring predefined hyperparameters, enabling user-adjustable exploration-exploitation trade-offs.
 Extensive experiments demonstrate that HERec consistently outperforms both Euclidean and hyperbolic baselines, achieving up to 5.49% improvement in utility metrics and 11.39% increase in diversity metrics, effectively mitigating information cocoons.
 1 INTRODUCTION
-As online platforms grow in size dramatically, users are overwhelmed by an influx of information, creating information cocoons that restrict their exposure to diverse and novel content (Kunaver & Požrl, 2017; Castells et al., 2021).
-Traditional recommendation approaches have primarily relied on collaborative filtering and interaction data (He et al., 2020; Koren et al., 2009), where users with similar historical preferences are expected to share future interests (Chen et al., 2021).
+As online platforms grow in size dramatically, users are overwhelmed by an influx of information, creating information cocoons that restrict their exposure to diverse and novel content [16, 4].
+Traditional recommendation approaches have primarily relied on collaborative filtering and interaction data [11, 14], where users with similar historical preferences are expected to share future interests [6].
 However, these methods face two critical challenges in improving user experience.
-First, existing models struggle to capture the underlying hierarchical structure inherent in user-item networks (Zheng et al., 2017; Unger & Tuzhilin, 2020), a critical factor for enhancing performance and personalization.
+First, existing models struggle to capture the underlying hierarchical structure inherent in user-item networks [48, 27], a critical factor for enhancing performance and personalization.
 Second, these models fail to effectively balance the exploration-exploitation trade-off.
-For instance, a jazz enthusiast may want to explore rock music, but current models primarily focus on exploitation—recommending only familiar genres—without adequately supporting exploration (as illustrated in Figure 1).
-Hyperbolic space has emerged as a promising approach for modeling hierarchical structures and power-law distributions (Krioukov et al., 2010; Yang et al., 2022b; Sun et al., 2021; Yang et al., 2022a), where its exponential volume growth naturally aligns with hierarchical, scale-free structures (Ganea et al., 2018; Chami et al., 2019; Liu et al., 2019; Zhang et al., 2021; Yang et al., 2023).
+For instance, a jazz enthusiast may want to explore rock music, but current models primarily focus on exploitation—recommending only familiar genres—without adequately supporting exploration.
+Hyperbolic space has emerged as a promising approach for modeling hierarchical structures and power-law distributions [15, 37, 25, 36], where its exponential volume growth naturally aligns with hierarchical, scale-free structures [10, 5, 18, 46, 38].
 In hyperbolic space, distances increase exponentially from the origin, allowing popular items or exploratory users to be positioned closer to the origin, while niche items or users with focused interests spread toward the boundaries.
-This spatial configuration enables hyperbolic graph neural network (GNN) models to perceive hierarchical relationships effectively, enhancing recommendation performance (Yang et al., 2022a; Sun et al., 2021; Yang et al., 2022b).
+This spatial configuration enables hyperbolic graph neural network (GNN) models to perceive hierarchical relationships effectively, enhancing recommendation performance [36, 25, 37].
 Despite these advantages, current hyperbolic recommender systems still face key limitations.
@@
 This mismatch leaves existing hyperbolic models vulnerable to noise in collaborative information, as they fail to capture semantic insights effectively.
-Second, hyperbolic recommender systems lack a principled approach to balancing exploration (introducing new content) and exploitation (focusing on existing preferences) (Wilson et al., 2021).
+Second, hyperbolic recommender systems lack a principled approach to balancing exploration (introducing new content) and exploitation (focusing on existing preferences) [32].
 This limitation reduces their ability to adapt to the exploration-exploitation trade-off, ultimately diminishing both the utility and diversity of recommendations.
-Proposed Methods.
+Proposed Method.
 To address these challenges, we propose HERec, a hyperbolic graph-llm framework designed to balance exploration and exploitation in recommender systems.
-Our approach introduces two key innovations: (1) a hierarchical-aware graph-LLM alignment mechanism, which jointly aligns textual descriptions with user-item collaborative information in hyperbolic space, supported by theoretical gradient analysis demonstrating its advantages, and (2) a novel hierarchical representation structure that enables user-adjustable exploration-exploitation trade-offs.
-Unlike conventional approaches that require manually defined hierarchies, we propose a hyperparameter-free clustering mechanism theoretically optimized by Dasgupta’s cost (Dasgupta, 2016).
+Our approach introduces two key innovations: (1) a hierarchical-aware graph-LLM alignment mechanism, which jointly aligns textual descriptions processed by LLMs with user-item collaborative information in hyperbolic space, supported by theoretical gradient analysis demonstrating its adaptiveness; (2) a novel hierarchical representation structure that enables user-adjustable exploration-exploitation trade-offs.
+Unlike conventional approaches that require manually defined hierarchies, we propose a hyperparameter-free clustering mechanism theoretically optimized by Dasgupta’s cost [7].
 This approach automatically discovers hierarchical structures without requiring predefined hyperparameters, allowing for a more adaptive and principled organization of user preferences.
 To validate the effectiveness of HERec, we conducted extensive experiments across multiple dimensions.
-Our model consistently outperformed baselines in both utility and diversity metrics, marking a significant achievement as it is the first model to excel in both aspects simultaneously.
-For a fair comparison, we enhanced Euclidean baselines with feature information, yet our model still demonstrated superior performance, validating the advantage of our designed alignment in hyperbolic space.
+Our model consistently outperformed baselines in both utility and diversity, marking a significant achievement as it is the first model to excel in both aspects simultaneously.
+For a fair comparison, we also enhanced Euclidean baselines with feature information, yet our model still demonstrated superior performance.
 Additionally, we conducted a detailed analysis on head and tail items, showing that HERec effectively boosts recommendations for tail items, thus enhancing diversity.
-An ablation study, along with a hierarchical representation structure analysis and decoder module, confirms HERec’s capability to leverage both semantic meaning and collaborative information, providing deeper insights into the advantages of our approach.
+An ablation study, along with a hierarchical representation structure analysis module, confirms HERec’s capability to leverage both semantic meaning and collaborative information, providing deeper insights into the advantages of our approach.
 Contributions.
@@
 The proposed graph-llm alignment framework unifies collaborative and semantic information in hyperbolic space, while a hierarchical representation structure organizes user preferences, allowing for dynamic adjustment between exploration and exploitation.
-Moreover, through extensive experiments, we demonstrate that HERec achieves state-of-the-art performance across both utility and diversity metrics, making it the first model to excel in both criteria simultaneously.
-To facilitate reproducibility and further research, we open-source our model implementation at: https://github.com/Martin-qyma/HERec.
-Figure 1: Exploration and exploitation in hyperbolic space.
+Moreover, through extensive experiments evaluating both utility and diversity metrics, we demonstrate that HERec achieves state-of-the-art performance across utility and diversity metrics, making it the first model to excel in both criteria simultaneously.
+We have made our model implementation available at: https://github.com/Martin-qyma/HERec.
 2 Related Work
@@
 Graph Collaborative Filtering has emerged as a dominant approach in recommender systems, with graph neural networks (GNNs) playing a central role.
-Unlike traditional collaborative filtering techniques, such as matrix factorization (Koren et al., 2009), GNNs capitalize on the relational structure in user-item interactions, enabling a more nuanced understanding of complex dependencies.
-Prominent graph-based methods, such as NGCF (Wang et al., 2019), SGL (Wu et al., 2021), and LightGCN (He et al., 2020), have been widely adopted and have consistently achieved state-of-the-art performance.
+Unlike traditional collaborative filtering techniques, such as matrix factorization [14], GNNs capitalize on the relational structure in user-item interactions, enabling a more nuanced understanding of complex dependencies.
+Prominent graph-based methods, such as NGCF [30], SGL [33], and LightGCN [11], have been widely adopted and have consistently achieved state-of-the-art performance.
 Nevertheless, these models are constrained to Euclidean space, which may fall short in capturing the underlying power-law distribution in user-item networks.
@@
 Large language models (LLMs) are increasingly used in recommender systems for their strong capability in understanding textual information.
-Some approaches leverage LLMs directly for downstream recommendation tasks (Bao et al., 2023; Li et al., 2023; Lyu et al., 2023; Zheng et al., 2024; Zhang et al., 2023), while others utilize LLMs as an auxiliary tool for collaborative filtering (Ren et al., 2024; Xi et al., 2024; Wei et al., 2024).
+Some approaches leverage LLMs directly for downstream recommendation tasks [1, 17, 19, 47, 45], while others utilize LLMs as an auxiliary tool for collaborative filtering [23, 34, 31].
 However, these approaches require substantial inference time and are computationally inefficient.
@@
 Hyperbolic Representation Learning.
-To address the limitations of Euclidean space in modeling hierarchical structures, researchers have investigated the use of hyperbolic space (Chami et al., 2019; Liu et al., 2019; Yang et al., 2022b; Zhang et al., 2021), where distances grow exponentially as embeddings move away from the origin.
+To address the limitations of Euclidean space in modeling hierarchical structures, researchers have investigated the use of hyperbolic space [5, 18, 37, 46], where distances grow exponentially as embeddings move away from the origin.
 This property makes hyperbolic geometry particularly well-suited for modeling user-item interactions by more accurately capturing latent hierarchies and dependencies.
-Consequently, hyperbolic representation learning has gained considerable attention in the recommender systems domain (Ying et al., 2018; Vinh Tran et al., 2020; Sun et al., 2021; Yang et al., 2022a), demonstrating improved performance over traditional Euclidean models.
+Consequently, hyperbolic representation learning has gained considerable attention in the recommender systems domain [41, 29, 25, 36], demonstrating improved performance over traditional Euclidean models.
 However, none of these hyperbolic models fully leverage the rich semantic information embedded in language-based datasets.
-While some approaches incorporate knowledge graphs as supplementary information (Tai et al., 2021; Du et al., 2022), these models primarily focus on refining hierarchical structures rather than achieving a deeper semantic understanding of each item or user.
+While some approaches incorporate knowledge graphs as supplementary information [26, 9], these models primarily focus on refining hierarchical structures rather than achieving a deeper semantic understanding of each item or user.
 3 PRELIMINARIES
-Embeddings learned from graph collaborative filtering are by default in the Euclidean space, which is unable to capture tree-like structures.
+Embeddings learned from graph collaborative filtering are by default in the Euclidean space, which is unable to capture hierarchical structures.
 To address the problem, we introduce how to derive these embeddings in a hyperbolic space.
 Riemannian Manifold.
-A Riemannian manifold $(\mathcal{M},g)$ is a smooth manifold that features an inner product $g_{x}:T_{x}\mathcal{M}\times T_{x}\mathcal{M}\rightarrow\mathbb{R}$ at each point $x\in\mathcal{M}$.
-This inner product, known as the Riemannian metric, defines geometric characteristics of the space such as angles and curve lengths.
-The curvature of Riemannian manifolds results in different types of geometries: elliptic geometry with positive curvature, Euclidean geometry with zero curvature, and hyperbolic geometry with negative curvature.
-In this work, we focus on hyperbolic geometry, characterized by negative curvature.
-Hyperbolic space can be described by several equivalent models, each offering unique properties while being mathematically equivalent.
-The Lorentz model, also referred to as the hyperboloid model, is one of the commonly used representations of hyperbolic space.
+A Riemannian manifold $(\mathcal{M},g)$ is a smooth manifold with an inner product $g_{x}$ at each point $x\in\mathcal{M}$, defining geometric properties such as angles and curve lengths.
+Its curvature determines the geometry: positive (elliptic), zero (Euclidean), or negative (hyperbolic).
+We focus on hyperbolic geometry, modeled using the Lorentz (hyperboloid) representation.
 Lorentz Model.
 An $n$-dimensional Lorentz manifold with negative curvature $-1/\kappa$ $(\kappa>0)$ is described as the Riemannian manifold $(\mathbb{H}^{n}_{\kappa},g_{\mathcal{L}})$.
-Here, $\mathbb{H}^{n}_{\kappa}=\{x\in\mathbb{R}^{n+1}:\langle x,x\rangle_{\mathcal{L}}=-\kappa,x_{0}>0\}$, $g_{\mathcal{L}}=\eta$ with $\eta=\mathbf{I}_{n}$ except $\eta_{0,0}=-1$, and $\langle\cdot,\cdot\rangle_{\mathcal{L}}$ represents the Lorentzian inner product, which is defined
+Here, $\mathbb{H}^{n}_{\kappa}=\{x\in\mathbb{R}^{n+1}:\langle x,x\rangle_{\mathcal{L}}=-\kappa,x_{0}>0\}$, $g_{\mathcal{L}}=\eta$ with $\eta=\mathbf{I}_{n}$ except $\eta_{0,0}=-1$, and $\langle\cdot,\cdot\rangle_{\mathcal{L}}$ represents the Lorentzian inner product, which is defined as:
 $\langle x,y\rangle_{\mathcal{L}}:=-x_{0}y_{0}+\sum_{i=1}^{n}x_{i}y_{i}$ (1)
-A tangent space is an $n$-dimensional vector space that approximates Euclidean space to hyperbolic space.
-Specifically, given a point $x\in\mathbb{H}_{\kappa}^{n}$, there exists a tangent space $\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}$ that approximates $\mathbb{H}_{\kappa}^{n}$:
+A tangent space is an $n$-dimensional vector space that locally approximates hyperbolic space.
+For a point $x\in\mathbb{H}_{\kappa}^{n}$, the tangent space $\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}$ is defined as:
 $\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}:=\{v\in\mathbb{R}^{n+1}:\langle v,x\rangle_{\mathcal{L}}=0\}$ (2)
-We also need to define transitions between $\mathbb{H}_{\kappa}^{n}$ and $\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}$, which give rise to the exponential and logarithmic maps.
-For $x\in\mathbb{H}_{\kappa}^{n}$ and $v\in\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}$ with $v\neq\mathbf{0}$ and $y\neq x$, there exists a unique geodesic $\gamma:[0,1]\to\mathbb{H}_{\kappa}^{n}$ such that $\gamma(0)=x$ and $\gamma^{\prime}(0)=v$.
-The exponential map $\exp_{x}:\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}\to\mathbb{H}_{\kappa}^{n}$ is given by $\exp_{x}(v)=\gamma(1)$:
+This space is orthogonal to $x$ under the Lorentzian inner product, providing accurate local approximation.
+Transitions between hyperbolic space and tangent space are facilitated by exponential and logarithmic maps, essential for gradient-based optimization.
+The exponential map $\exp_{x}:\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}\to\mathbb{H}_{\kappa}^{n}$ projects a tangent vector onto hyperbolic space along a geodesic:
 $\exp_{x}^{\kappa}(v)=\cosh\left(\frac{\|v\|_{\mathcal{L}}}{\sqrt{\kappa}}\right)x+\sqrt{\kappa}\sinh\left(\frac{\|v\|_{\mathcal{L}}}{\sqrt{\kappa}}\right)\frac{v}{\|v\|_{\mathcal{L}}}$ (3)
-where $\|v\|_{\mathcal{L}}=\sqrt{\langle v,v\rangle_{\mathcal{L}}}$ denotes the Lorentzian norm of $v$.
-The logarithmic map $\log_{x}$ is the inverse of the exponential map $\exp_{x}$, and the distance between $x$ and $y$ in $\mathbb{H}^{n}_{\kappa}$ is:
+where $\|v\|_{\mathcal{L}}=\sqrt{\langle v,v\rangle_{\mathcal{L}}}$ is the Lorentzian norm.
+The logarithmic map $\log_{x}$ is the inverse of $\exp_{x}$ and retrieves the tangent vector for a point $y\in\mathbb{H}_{\kappa}^{n}$:
 $\log_{x}^{\kappa}(y)=d_{\mathcal{L}}^{\kappa}(x,y)\frac{y+\frac{1}{\kappa}\langle x,y\rangle_{\mathcal{L}}x}{\|y+\frac{1}{\kappa}\langle x,y\rangle_{\mathcal{L}}x\|}$ (4)
@@
 $d_{\mathcal{H}}^{\kappa}(x,y)=\sqrt{\kappa}\,\text{arcosh}\left(-\langle x,y\rangle_{\mathcal{L}}/\kappa\right)$ (5)
-We choose $\mathbf{o}:=\{\sqrt{\kappa},0,\ldots,0\}\in\mathbb{H}^{n}_{\kappa}$ as the reference point for all the operations, making the transitions reproducible.
-In this work, we set $\kappa=1$, resulting in a curvature of $-1$.
-Therefore, in the following discussions, we would omit $\kappa$ for simplicity.
+We use $\mathbf{o}:=\{\sqrt{\kappa},0,\ldots,0\}\in\mathbb{H}^{n}_{\kappa}$ as the reference point.
+For simplicity, we set $\kappa=1$, giving a curvature of $-1$, and omit $\kappa$ in subsequent discussions.
 4 METHODOLOGY
@@
 4.1 Hyperbolic Graph Collaborative Filtering
-Similar to Euclidean space, hyperbolic graph collaborative filtering captures user-item dependencies via message passing mechanism.
-In our method, we use the Lorentz representation for both users and items.
+Similar to Euclidean space, hyperbolic graph collaborative filtering captures user-item dependencies via message passing mechanism. We apply the Lorentz representation for both users and items.
 Hyperbolic embedding initialization.
 As discussed in section 3, we fix the origin $\mathbf{o}:=\{\sqrt{\kappa},0,\ldots,0\}\in\mathbb{H}^{n}_{\kappa}$ and use it as the reference point.
-We adopt the Gaussian sampling method for the initialization of embeddings in Euclidean space, after that, by injecting them into the tangent space of the reference point, we derive the initial hyperbolic embeddings of both users and items:
+Embeddings are initialized in Euclidean space using Gaussian sampling and projected onto the tangent space of $\mathbf{o}$ to derive their hyperbolic representations:
 $z_{u}^{0}=(0,e_{u}),\quad\qquad z_{i}^{0}$ $=(0,e_{i})$ (6)
 $h_{u}^{0}=\exp_{\mathbf{o}}(z_{u}^{0}),\quad\quad h_{i}^{0}$ $=\exp_{\mathbf{o}}(z_{i}^{0})$
-where $e_{u},e_{i}$ are Gaussian distributed initialized Euclidean embeddings, and $z_{u}^{0},z_{i}^{0}$ are corresponding embeddings in the tangent space of the origin point.
-$h_{u}^{0},h_{i}^{0}$ therefore stands for the resulting initialized hyperbolic embeddings.
+where $e_{u},e_{i}$ are Gaussian-initialized Euclidean embeddings, and $z_{u}^{0},z_{i}^{0}$ are their tangent space projections. $h_{u}^{0},h_{i}^{0}$ denote the resulting hyperbolic embeddings.
 Hyperbolic Message Passing.
-Message passing between each node is realized through hyperbolic neighborhood aggregation, which is used to extract explicit user-item interaction.
-Given user $u$, item $i$, as well as their neighborhood $\mathcal{N}_{u},\mathcal{N}_{i}$, their embeddings at layer $l$ in the tangent space are updated as follows:
-$z_{u}^{l}=z_{u}^{l-1}+\sum_{i\in\mathcal{N}_{u}}\frac{1}{|\mathcal{N}_{u}|}z_{i}^{l-1},\ z_{i}^{l}=z_{i}^{l-1}+\sum_{u\in\mathcal{N}_{i}}\frac{1}{|\mathcal{N}_{i}|}z_{u}^{l-1}$ (7)
-where $|\mathcal{N}_{u}|$ and $|\mathcal{N}_{i}|$ represent the number of neighbors of $u$ and $i$, respectively. The final tangent space embedding is derived by summing up embeddings at each layer:
+User-item interactions are captured via hyperbolic neighborhood aggregation.
+For user $u$, item $i$, and their neighborhoods $\mathcal{N}_{u},\mathcal{N}_{i}$, tangent space embeddings at layer $l$ are updated as:
+$z_{u}^{l}=z_{u}^{l-1}+\sum_{i\in\mathcal{N}_{u}}\frac{1}{|\mathcal{N}_{u}|}z_{i}^{l-1},\quad z_{i}^{l}=z_{i}^{l-1}+\sum_{u\in\mathcal{N}_{i}}\frac{1}{|\mathcal{N}_{i}|}z_{u}^{l-1}$ (7)
+where $|\mathcal{N}_{u}|$ and $|\mathcal{N}_{i}|$ are the neighborhood sizes. The final tangent embeddings are obtained by summing over all layers:
 $z_{u}=\sum_{l}z_{u}^{l},\quad\quad z_{i}=\sum_{l}z_{i}^{l}$ (8)
-Since $z_{u}$ and $z_{i}$ are on the tangent space, we still need to project it into hyperbolic space to derive the hyperbolic embeddings:
+Since $z_{u}$ and $z_{i}$ are on the tangent space, we still need to project them into hyperbolic space to derive the hyperbolic embeddings:
 $h_{u}=\exp_{\mathbf{o}}(z_{u}),\quad\quad h_{i}=\exp_{\mathbf{o}}(z_{i})$ (9)
@@
 Hyperbolic Margin Ranking and Informative Mechanism.
-To address the limitations of Euclidean-based approaches, we adopt the hyperbolic margin ranking loss (Sun et al., 2021), an extension of Bayesian Personalized Ranking (BPR) (Rendle et al., 2012).
+To address the limitations of Euclidean-based approaches, we adopt the hyperbolic margin ranking loss [25], an extension of Bayesian Personalized Ranking (BPR) [24].
 This loss pulls user-item pairs with observed interactions closer while pushing negative pairs apart, preventing the collapse seen in Euclidean space by incorporating a margin:
@@
 where $d_{\mathcal{H}}$ represents hyperbolic distance, $m$ is the margin, $(u,i)$ is an observed interaction, and $(u,j)$ is a sampled negative pair.
-To further refine this pull-push process, we integrate a geometric-aware mechanism (Yang et al., 2022a) combining Hyperbolic-aware Margin Learning (HAML) and Hyperbolic Informative Negative Sampling (HINS).
+To further refine this pull-push process, we integrate a geometric-aware mechanism [36] combining Hyperbolic-aware Margin Learning (HAML) and Hyperbolic Informative Negative Sampling (HINS).
 HAML dynamically adjusts the margin based on hyperbolic geometry, assigning a larger margin to pairs near the origin:
@@
 $\delta$ $=\frac{d_{\mathcal{H}}^{2}(e_{u},\mathbf{o})+d_{\mathcal{H}}^{2}(e_{i},\mathbf{o})-d_{\mathcal{H}}^{2}(e_{u},e_{i})}{e_{u,0}\cdot e_{i,0}}$
-where $e_{u,0}$ and $e_{i,0}$ are the zeroth coordinate elements of $e_{u}$ and $e_{i}$, respectively, corresponding to the embeddings norms in the hyperbolic space.
-HINS enhances negative sampling by prioritizing candidates close to the positive item, iteratively identifying the most informative negative samples.
-This method leverages hyperbolic geometry to effectively model head and tail items, improving utility performance in recommendations.
+where $e_{u,0}$ and $e_{i,0}$ are the zeroth coordinates of $e_{u}$ and $e_{i}$, representing their norms.
+HINS enhances negative sampling by prioritizing candidates near the positive item, leveraging hyperbolic geometry to better model head and tail items, ultimately improving recommendation performance.
 4.2 Semantic Representations Generation
-In this subsection, we introduce a paradigm for distilling useful information from raw text descriptions of items and users into dense semantic embeddings, which will be used in our alignment framework to derive the final representation.
-Inspired by prior research (Ma et al., 2024; Ren et al., 2024) that leverages the power of large language models (LLMs) to extract useful user/item descriptions from raw, noisy textual data, we adopt a similar framework to generate user/item profiles with rich textual descriptions: $\mathcal{P}_{i}=\textit{LLMs}(\mathcal{S}_{i},\mathcal{T}_{i})$, where $\mathcal{S}_{i}$ represents system prompts to the LLMs, $\mathcal{T}_{i}$ denotes raw textual information, and $\mathcal{P}_{i}$ is the resulting user/item profile, consisting of sentence-level descriptions.
-While we’ve derived textual information for users/items, it cannot be directly fused with the embeddings generated from hyperbolic collaborative filtering due to discrepancies in embedding geometry.
-To integrate both types of knowledge, we encode the textual information in embeddings using pre-trained text encoder models: $e_{i}=\textit{Encoder}(\mathcal{P}_{i})$.
-While LLM embeddings could be an alternative, they often introduce higher computational overhead and may not be optimized for alignment with hyperbolic representations, making encoders a more practical choice.
+We propose a paradigm to distill useful information from raw text descriptions of items and users into dense semantic embeddings for our alignment framework.
+Building on prior work [20, 23] that leverages large language models (LLMs) to extract meaningful user/item descriptions from noisy textual data, we generate rich user/item profiles as $\mathcal{P}_{i}=\textit{LLMs}(\mathcal{S}_{i},\mathcal{T}_{i})$, where $\mathcal{S}_{i}$ are system prompts, $\mathcal{T}_{i}$ is raw textual input, and $\mathcal{P}_{i}$ contains sentence-level descriptions.
+To integrate these profiles with hyperbolic collaborative filtering embeddings, we encode the textual information into embeddings using pre-trained text encoders: $e_{i}=\textit{Encoder}(\mathcal{P}_{i})$.
+While LLM embeddings are an alternative, they often incur higher computational overhead and are less optimized for alignment with hyperbolic representations, making pre-trained encoders a more practical choice.
 4.3 Hyperbolic Alignment
-We propose maximizing mutual information between hyperbolic embeddings (collaborative information) and semantic embeddings (textual information) to create unified embeddings that integrate both knowledge types.
-To achieve this, we introduce a hyperbolic alignment loss, which reduces noise in representations and improves tail item recommendations by aligning information in the same hyperbolic space.
+To maximize mutual information between hyperbolic embeddings (collaborative information) and semantic embeddings (textual information), we introduce a hyperbolic alignment framework to create unified representations that integrate both knowledge types.
+To achieve this, we develop a hyperbolic alignment loss, which reduces noise in representations and improves tail item recommendations by aligning information in the same hyperbolic space.
 A theoretical advantage of hyperbolic alignment over Euclidean alignment is discussed in the next subsection.
@@
 To accomplish this, we first apply a multi-layer perceptron (MLP) to the semantic embeddings to ensure dimensional alignment.
-We then project the semantic embeddings into hyperbolic space so they can align with the hyperbolic collaborative embeddings:
-$s_{i}=\exp_{x}(e_{i}),\quad s^{\prime}_{i}=\textit{MLP}(s_{i})$ (13)
-where $\exp_{x}$ is defined as in Equation 3, $e_{i}$ represents the semantic embeddings from the encoder, $s_{i}$ is the dimension-aligned semantic embedding, and $s^{\prime}_{i}$ denotes the final projected semantic embedding in hyperbolic space.
-The hyperbolic alignment loss is then calculated as:
-$\ell_{align}(i)=d_{\mathcal{H}}^{2}(h_{i},s^{\prime}_{i})$ (14)
-where $d_{\mathcal{H}}$ represents the hyperbolic distance, and $h_{i}$ is the hyperbolic embedding with collaborative information.
+We then project the semantic embeddings into hyperbolic space so they can align with the hyperbolic collaborative embeddings.
+The hyperbolic alignment loss is calculated as:
+$\ell_{align}(i)=d_{\mathcal{H}}^{2}(h_{i},s_{i})$ (13)
+where $d_{\mathcal{H}}$ represents the hyperbolic distance, $h_{i}$ is the hyperbolic embedding with collaborative information and $s_{i}$ denotes the projected semantic embedding in hyperbolic space.
 4.4 Gradient Analysis of Hyperbolic Alignment
-To further explore the impact of aligning semantic embeddings within the hyperbolic model, we analyze the gradient behavior with respect to nodes of varying norms in Proposition 4.1, proof can be found in Appendix A.
+To further explore the impact of aligning semantic embeddings within the hyperbolic model, we analyze the gradient behavior with respect to nodes of varying norms in Proposition 1, detailed proof with analysis can be found in Appendix A.
 This provides insight into the hierarchical preservation properties of the hyperbolic space, which are absent in Euclidean space.
-Proposition 4.1.
+Proposition 1.
 Let $\mathbf{x},\mathbf{y}$ denote embeddings in hyperbolic space $\mathcal{H}^{n}$ with corresponding unit direction vectors in space-like dimension $\hat{\mathbf{x}},\hat{\mathbf{y}}$, and let $\theta$ be the angle between them.
@@
 Meanwhile, nodes with smaller norms (abstract preferences, positional changes have less impact on distances) adjust more, refining global relationships and effectively capturing the hierarchical structure.
+This inherent ability of hyperbolic space to modulate learning intensity based on an embedding’s position within the hierarchy offers a distinct advantage over the fixed-update approach of Euclidean geometry, ultimately facilitating more faithful and structured representations.
 4.5 Hierarchical Representation Structure
 In this subsection, we design a mechanism that allows users to balance the exploration-exploitation trade-off by constructing a hierarchical representation structure based on the learned hyperbolic embeddings.
+A detailed complexity analysis is provided in Appendix C.2.
 Hierarchy Tree.
@@
 However, the number of cluster nodes in each layer is not predefined and must be determined.
-Inspired by Dasgupta’s cost (Dasgupta, 2016), which encourages pushing edge cuts as far down the tree as possible, the optimal tree is required to be binary.
+Inspired by Dasgupta’s cost [7], which encourages pushing edge cuts as far down the tree as possible, the optimal tree is required to be binary.
 Therefore, we fix the proportion of cluster nodes between layers as $k=2$.
-The details of the algorithm are presented in Algorithm 1.
-Algorithm 1 Hyperbolic Hierarchical Clustering
-  Input: Hyperbolic embeddings $X$, layer proportion $k$
-  Set the maximum number of layers $L=\log_{k}(|X|)$
-  Initialize the points of bottom layer $D_{L}=X$
-  for $l=L$ to $1$ (bottom to top) do
-     Randomly initialize cluster centroids in the hyperbolic space $M_{l}=\{\mu_{l}\ |\ \mu_{l}\in\mathbb{H}\}$, $|M_{l}|=|D_{l}|/k$
-     repeat
-        for each point $d_{l,i}\in D_{l}$ do
-           Assign $d_{i}$ to the cluster with the nearest centroid:
-           $c_{l,i}=\arg\min_{\mu\in M_{l}}d_{\mathcal{H}}(d_{i},\mu)$
-        end for
-        for each cluster $c_{l,i}\in C_{l}$ do
-           Update the centroid $\mu_{l,i}$ as the hyperbolic mean $m_{\mathcal{H}}$ of all points assigned to cluster $c_{l,i}$:
-           $\mu_{l,i}=m_{\mathcal{H}}(\{d\ |\ d\in C_{l,i}\})$
-        end for
-     until Centroids $M_{l}$ do not change (convergence)
-     Initialize the points of the upper layer $D_{l-1}=M_{l}$
-  end for
+Details of the algorithm are presented in Appendix C.1.
 Exploration-Exploitation Balance.
@@
 In that case, half of the original recommendations are retained, and the other half is replaced with items sampled from its ancestor cluster at layer $5$ of the hierarchy tree.
-Complexity Analysis.
-To demonstrate the scalability and efficiency of our approach, we analyze the complexity of constructing and applying the hierarchical representation structure.
-Let $N$ denote the total number of users and items, and let $k$ be the number of recommendations per user.
-The hierarchy tree is built using iterative $k$-means clustering in hyperbolic space, where cluster nodes are assigned only to real node embeddings for efficiency.
-First, computing and storing pairwise distances requires $O(N^{2}/2)$ operations.
-Each $k$-means iteration has complexity $O(i)$, and with $O(\log N)$ hierarchical levels, the total complexity of constructing the hierarchy tree is:$O\left(\frac{N^{2}}{2}+i\log N\right)$.
-For recommendation retrieval, finding the top-$k$ recommendations involves nearest neighbor search in hyperbolic space, which runs in $O(k\log N)$.
-Adjusting recommendations by retrieving ancestor clusters at layer $l$ requires only tree traversal, which runs in $O(k)$.
-Overall, our hierarchical framework ensures efficient scaling while maintaining flexibility in recommendation adjustments.
-This design enables HERec to dynamically balance exploration and exploitation while ensuring scalability for large-scale recommender systems.
 5 EXPERIMENTS
@@
 Ours 0.1002 0.0772 0.1598 0.0992 0.3185 12.466 0.8598 0.3543 12.769 0.8728
-Table 2: Comparison with feature-enhanced Euclidean baselines, with the best-performing results highlighted in bold.
-Data Amazon-books Yelp Google-reviews
-Metrics R@20 N@20 Div@20 H@20 EPC@20 R@20 N@20 Div@20 H@20 EPC@20 R@20 N@20 Div@20 H@20 EPC@20
-LightGCN+ 0.1466 0.0901 0.4349 10.3915 0.7802 0.1156 0.0733 0.2697 9.6903 0.7526 0.1438 0.0940 0.2707 10.8664 0.7801
-LightGCL+ 0.1545 0.0992 0.5064 11.9761 0.8728 0.1131 0.0739 0.2759 11.8617 0.8581 0.1443 0.0946 0.2744 12.4700 0.8776
-SimGCL+ 0.1494 0.0926 0.5163 11.4751 0.8342 0.1195 0.0752 0.3152 9.9952 0.7811 0.1504 0.0982 0.3030 11.2111 0.7953
-HCCF+ 0.1525 0.0940 0.4316 11.7984 0.8616 0.1170 0.0729 0.2884 11.6906 0.8577 0.1438 0.0941 0.2524 11.5433 0.8406
-Ours 0.1682 0.1069 0.5754 12.2543 0.8758 0.1224 0.0762 0.3961 12.3290 0.8773 0.1598 0.0992 0.3543 12.7690 0.8728
-Table 3: Statistics of the experimental data.
-Dataset #User #Item #Interactions Density
-All H20(%) T80(%)
-[…]
-[…]
-[…]
+5.1 Experimental Settings
+Datasets.
+To evaluate our proposed model, we utilize three widely-used public datasets: Amazon-books, which contains user purchase behaviors within the book category on Amazon; Yelp, which records customer ratings and reviews for restaurants on Yelp; and Google-reviews, which includes user reviews and business metadata from Google Maps.
+Detailed statistics, including the distribution of head and tail items, are provided in Appendix B.
+The high proportion of tail users (T80) highlights the prevalence of long-tail distributions, supporting the assumption of a power-law distribution in real-world data.
+More implementation details can be found in Appendix B.1.
+Evaluation Metrics.
+We assess our recommended items using both utility and diversity metrics, demonstrating that our model achieves precise recommendations for users’ preference while enhancing diversity in the recommendation.
+For utility, we adopt two widely used ranking metrics: Recall and NDCG, which measure the effectiveness of the model.
+To evaluate diversity, we employ three metrics: Distance Diversity (Div) [49, 43], Shannon Entropy (H) [21], and Expected Popularity Complement (EPC) [28].
+For more details, please refer to Appendix B.2.
+Baselines. We compare our model with state-of-the-art Euclidean and Hyperbolic baselines:
+  • LightGCN [11]: A lightweight graph convolution neural network model for recommendation by removing redundant neural modules in graph convolution.
+  • TAG-CF [12]: A test-time augmentation framework that enhances recommendation by performing a single message-passing step only at inference time.
+  • LightGCL [3]: A contrastive learning model that uses singular value decomposition (SVD) for robust graph augmentation without data augmentation.
+  • SimGCL [42]: A contrastive learning model that replaces graph augmentations in contrastive learning by introducing random noise during initialization.
+  • HCCF [35]: A collaborative filtering framework that integrates hypergraph and contrastive learning, capturing both local and global user dependencies.
+  • HGCF [25]: A hyperbolic GNN model that combines hypergraph learning with collaborative filtering using hyperbolic margin ranking loss.
+  • HICF [36]: A hyperbolic recommendation model that enhances head and tail item performance with hyperbolic geometry, improving utility for long-tail items.
+5.2 Performance Comparison
+Overall Comparison.
+The performance comparison across all models is presented in Table 1, with statistical significance testing in Appendix D.2.
+Our HERec consistently outperforms the baselines in both utility and diversity metrics, achieving up to $5.49\%$ improvement in utility metrics and $11.39\%$ increase in diversity metrics.
+Notably, most baselines struggle to achieve strong performance in both utility and diversity simultaneously.
+For instance, while HICF performs second-best in utility metrics on the Amazon-books dataset, it lacks the same advantage in diversity metrics, where several baselines achieve better results.
+This highlights the superiority of our model, which achieves state-of-the-art performance in both utility and diversity: a feat not previously attained by any baseline.
+Baselines with Feature Enhancement.
+To ensure a fair comparison, we further enhance Euclidean baselines by integrating feature information.
+Specifically, we align semantic embeddings with the original Euclidean embeddings following the structure in Section 4.2, but in Euclidean space.
+The results are presented in Appendix D.1, with enhanced baselines denoted by “+”.
+The data shows that our HERec still surpasses all baselines in both utility and diversity.
+Interestingly, while the feature-enhanced versions of the Euclidean baselines generally improve in utility metrics, they do not exhibit a stable improvement in diversity metrics.
+This aligns with our theoretical analysis in Section 4.4, which demonstrates the advantages of our alignment in the hyperbolic space.
+Profile Decoder.
+To validate the effectiveness of our alignment framework in capturing semantics, we designed a profile decoder module to reconstruct user and item profiles from the learned hyperbolic embeddings.
+The results demonstrate that profiles reconstructed from our model’s embeddings exhibit higher alignment with the originals, confirming the model’s ability to effectively capture underlying semantic meaning.
+Detailed methodology and experimental results are provided in Appendix E.
+Performance on Head and Tail Items.
+To underscore the effectiveness of semantic alignment, particularly for cold-start (tail) items, we conducted experiments measuring utility performance on both head and tail items.
+In our context, head items (H20) refer to the top $20\%$ of items with the highest number of interactions in the training dataset, representing the most popular items.
+The remaining items are categorized as tail items (T80), which have limited interaction history, representing less frequently engaged items.
+For simplicity, we compare our model with all hyperbolic baselines and a representative Euclidean baseline (LightGCN).
+The results, displayed in Table 2, reveal significant improvements, especially for tail items.
+Compared with hyperbolic baselines without semantic alignment, our model shows a notably larger performance boost on tail items than on head items.
+This finding indicates that semantic alignment particularly enhances the recommendation of cold-start items, effectively mitigating the cold-start problem for items with limited or no prior interactions.
 Table 2: Performance on H20 (head) and T80 (tail) items, with the top hyperbolic results highlighted in bold.
-$\Delta_{\mathcal{H}}(\%)$ represents the improvement over the second-best hyperbolic baseline, which is underlined.
+$\Delta_{\mathcal{H}}(\%)$ represents the improvement over the second-best hyperbolic baseline (underlined).
 Datasets Amazon-books Yelp Google-reviews
@@
 $\Delta_{\mathcal{H}}(\%)$ +3.59 +8.58 +2.38 +9.71 +1.89 +12.11 +0.36 +13.33 +0.70 +14.18 +1.21 +14.42
-5.1 Experimental Settings
-Datasets.
-To evaluate our proposed model, we utilize three widely-used public datasets: Amazon-books, which contains user purchase behaviors within the book category on Amazon; Yelp, which records customer ratings and reviews for restaurants on Yelp; and Google-reviews, which includes user reviews and business metadata from Google Maps.
-Detailed statistics, including the distribution of head and tail items, are provided in Table 3.
-The high proportion of tail users (T80) highlights the prevalence of long-tail distributions, supporting the assumption of a power-law distribution in real-world data.
-Evaluation Metrics.
-We assess our recommended items using both utility and diversity metrics, demonstrating that our model achieves precise recommendations for users’ preference while enhancing diversity in the recommendation.
-For utility, we adopt two widely used ranking metrics: Recall and NDCG, which measure the effectiveness of the model.
-To evaluate diversity, we employ three metrics: Distance Diversity (Div) (Ziegler et al., 2005; Zhang & Hurley, 2008), Shannon Entropy (H) (Mehta et al., 2012), and Expected Popularity Complement (EPC) (Vargas & Castells, 2011).
-Distance Diversity (Div) quantifies the variation among recommended items, encouraging diversity within the recommendation:
-$Div(u)=\frac{\sum_{i\in R_{u}}\sum_{j\in R_{u},j\neq i}d(i,j)}{\text{Number of}(i,j)\text{ pairs}}$ (16)
-where $R_{u}$ is the recommendation list for user $u$, and $d(i,j)$ represents the distance between items $i$ and $j$.
-Number of $(i,j)$ pairs is calculated as $\binom{|R_{u}|}{2}=\frac{|R_{u}|(|R_{u}|-1)}{2}$.
-For fair comparison, we compute distances within the same representation space across models.
-Shannon Entropy (H) quantifies the unpredictability within the recommendations, with higher entropy signifying greater diversity:
-$H=-\sum_{i\in\text{set}(\sum_{u}R_{u})}p(i)\log p(i)$ (17)
-where $\sum_{u}R_{u}$ denotes all recommended items across users (including duplicates), and $p(i)=\text{count}(i)/\sum_{i\in\text{set}(\sum_{u}R_{u})}\text{count}(i)$ is the proportion of item $i$ among all recommended items.
-Expected Popularity Complement (EPC) assesses the extent of bias towards popular items by promoting niche recommendations:
-$EPC=1-\frac{1}{|\text{set}(\sum R_{u})|}\sum_{i\in\text{set}(\sum_{u}R_{u})}\frac{\text{pop}(i)}{\max_{i}(\text{pop}(i))}$ (18)
-where $N_{u}$ is the set of items user $u$ has interacted with, and $\text{pop}(i)=\text{count}(i)/\sum_{i\in\text{set}(\sum_{u}N_{u})}\text{count}(i)$ denotes item $i$’s popularity.
-Baselines. We compare our model with state-of-the-art baselines in both Euclidean and Hyperbolic spaces:
-  • LightGCN (He et al., 2020): A lightweight graph convolution neural network model for recommendation by removing redundant neural modules in graph convolution.
-  • TAG-CF (Ju et al., 2024): A test-time augmentation framework that enhances recommendation by performing a single message-passing step only at inference time.
-  • LightGCL (Cai et al., 2023): A contrastive learning model that uses singular value decomposition (SVD) for robust graph augmentation without data augmentation.
-  • SimGCL (Yu et al., 2022): A contrastive learning model that replaces graph augmentations in contrastive learning by introducing random noise during initialization.
-  • HCCF (Xia et al., 2022): A collaborative filtering framework that integrates hypergraph and contrastive learning, capturing both local and global user dependencies.
-  • HGCF (Sun et al., 2021): A hyperbolic GNN model that combines hypergraph learning with collaborative filtering using hyperbolic margin ranking loss.
-  • HICF (Yang et al., 2022a): A hyperbolic recommendation model that enhances head and tail item performance with hyperbolic geometry, improving utility for long-tail items.
-Training Details.
-We set embedding dimensions to 50 and a training batch size of 1024.
-Embeddings in hyperbolic space are optimized using Riemannian SGD (RSGD) (Bécigneul & Ganea, 2018; Sun et al., 2021), while the MLP adapter is optimized with Adam (Kingma, 2014).
-Training incorporates early stopping technique based on Recall@10 with 10 patience steps.
-We generate user and item profiles using GPT-3.5-turbo and encode them with text-embedding-ada-002 (Neelakantan et al., 2022) for Amazon-books and Yelp, and BERT (Devlin, 2018) for Google-reviews.
-All experiments are conducted on NVIDIA A40/A100 GPUs with 40GB memory.
 Figure 2: Ablation study on model variants.
-5.2 Performance Comparison
-Overall Comparison.
-The performance comparison across all models is presented in Table 1.
-Our HERec consistently outperforms the baselines in both utility and diversity metrics, achieving up to $5.49\%$ improvement in utility metrics and $11.39\%$ increase in diversity metrics.
-Notably, most baselines struggle to achieve strong performance in both utility and diversity simultaneously.
-For instance, while HICF performs second-best in utility metrics on the Amazon-books dataset, it lacks the same advantage in diversity metrics, where several baselines achieve better results.
-This highlights the superiority of our model, which achieves state-of-the-art performance in both utility and diversity: a feat not previously attained by any baseline.
-Baselines with Feature Enhancement.
-To ensure a fair comparison, we further enhance Euclidean baselines by integrating feature information.
-Specifically, we align semantic embeddings with the original Euclidean embeddings following the structure in Section 4.2, but in Euclidean space.
-The results are presented in Table 2, with enhanced baselines denoted by “+”.
-The data shows that our HERec still surpasses all baselines in both utility and diversity.
-Interestingly, while the feature-enhanced versions of the Euclidean baselines generally improve in utility metrics, they do not exhibit a stable improvement in diversity metrics.
-This aligns with our theoretical analysis in Section 4.4, which demonstrates the advantages of our alignment in the hyperbolic space.
-Additionally, to validate the effectiveness of our alignment framework in capturing semantics, we designed a profile decoder module to reconstruct user and item profiles from the learned hyperbolic embeddings.
-The results demonstrate that profiles reconstructed from our model’s embeddings exhibit higher alignment with the originals, confirming the model’s ability to effectively capture underlying semantic meaning.
-Detailed methodology and experimental results are provided in Appendix B.
+Figure 3: Layer-wise norms.
 Figure 4: Performance analysis across hierarchical structure layers.
-Figure 5: Layer-wise norms within the hierarchy structure.
-Performance on Head and Tail Items.
-To underscore the effectiveness of semantic alignment, particularly for cold-start (tail) items, we conducted experiments measuring utility performance on both head and tail items.
-In our context, head items (H20) refer to the top $20\%$ of items with the highest number of interactions in the training dataset, representing the most popular items.
-The remaining items are categorized as tail items (T80), which have limited interaction history, representing less frequently engaged items.
-For simplicity, we compare our model with all hyperbolic baselines and a representative Euclidean baseline (LightGCN).
-The results, displayed in Table 4, reveal significant improvements, especially for tail items.
-Compared with hyperbolic baselines without semantic alignment, our model shows a notably larger performance boost on tail items than on head items.
-This finding indicates that semantic alignment particularly enhances the recommendation of cold-start items, effectively mitigating the cold-start problem for items with limited or no prior interactions.
 5.3 Ablation Study
@@
 Models lacking both margin ranking and semantic alignment perform the worst, while removing either module results in decreased utility, emphasizing the effectiveness of integration.
+Recent work [40, 39] has proposed performing message passing directly in hyperbolic space as an alternative to using the tangent space.
+We implemented this approach in place of our original hyperbolic graph collaborative filtering method.
+As shown in Appendix D.3, this alternative achieves comparable performance, making it a viable alternative.
 5.4 Hierarchical Representation Structure Analysis
@@
 This structure provides users with a flexible mechanism to adjust recommendations from niche, relevant items to broader, exploratory sets based on their preferences.
-A case study is provided in Appendix C.
-Additionally, to validate the hierarchical structure’s capability to capture varying degrees of group preferences, we analyzed the average norm of node embeddings across layers (Figure 5).
+A case study is provided in Appendix C.3.
+To further validate the hierarchical structure’s capability to capture varying degrees of group preferences, we analyzed the average norm of node embeddings across layers (Figure 3).
 The results reveal a consistent increase at larger layers (farther from the center in hyperbolic space).
 As previously mentioned, hyperbolic geometry ensures that popular items or exploratory users are positioned closer to the origin, while niche items or users are positioned toward the boundaries.
-This finding aligns seamlessly with the properties of hyperbolic space, validating the intended design of the hierarchical structure: higher-level (smaller-layer) nodes effectively encapsulate abstract and aggregated preferences, whereas lower-level (larger-layer) nodes focus on fine-grained, personalized preferences.
+This observation aligns with the theoretical properties of hyperbolic space and supports our hierarchical design: higher-level (smaller-layer) nodes effectively encapsulate abstract and aggregated preferences, whereas lower-level (larger-layer) nodes focus on fine-grained, personalized preferences.
 6 CONCLUSION
 In this work, we present HERec, a hyperbolic graph-llm framework that effectively balances exploration and exploitation in recommender systems.
-Our approach integrates semantic and collaborative information within hyperbolic space, leveraging the unique properties of hyperbolic geometry.
-Furthermore, we introduce a personalized mechanism for balancing the exploration-exploitation trade-off using a hierarchical representation structure, empowering users to adjust recommendation preferences flexibly.
-Extensive experiments demonstrate that our model achieves state-of-the-art performance, achieving up to a $5.49\%$ improvement in utility metrics and an $11.39\%$ increase in diversity metrics.
+Our approach integrates semantic and collaborative information within hyperbolic space, leveraging the unique properties of hyperbolic geometry, supported by rigorous mathematical proof.
+Furthermore, we introduce a personalized mechanism for balance the exploration-exploitation trade-off using a hierarchical representation structure, empowering users to adjust recommendation preferences flexibly.
+Extensive experiments demonstrate that HERec achieves state-of-the-art performance, delivering up to a $5.49\%$ improvement in utility metrics and an $11.39\%$ increase in diversity metrics.
 Remarkably, HERec is the first framework to excel in both utility and diversity simultaneously, marking a significant advancement in the field.
-ACKNOWLEGDEMENTS
-This work was partially supported by the National Science Foundation (NSF) IIS Div Of Information & Intelligent Systems 2403317.
-We also gratefully acknowledge partial support from an Amazon Research Award, funding generously provided by Snap Research, and resources made available by the Yale Office of the Provost.
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
 Appendix A Proof
-In this section, we provide a rigorous mathematical proof of Proposition 4.1, demonstrating the unique advantage of the aligning mechanism in the hyperbolic space.
+In this section, we provide a rigorous mathematical proof of Proposition 1, demonstrating the unique advantage of the aligning mechanism in the hyperbolic space.
 A.1 Gradient Analysis in Hyperbolic Space.
@@
 This fundamental difference in gradient dynamics explains hyperbolic space’s superiority for hierarchical representation learning.
-Appendix B Profile Decoder
-Figure 6: A depiction of model prompt instruction.
-To further evaluate the effectiveness of hyperbolic alignment and the expressiveness of our learned hyperbolic embeddings, we introduce a profile decoder designed to reconstruct the original profile from these embeddings, thereby broadening their application beyond recommendation tasks.
-Our method begins by mapping the hyperbolic embeddings back into Euclidean space to facilitate alignment with traditional network layers.
-These Euclidean-mapped embeddings are then processed through a mixture of experts (MoE) module.
-The MoE serves to refine the embeddings, ensuring consistency in both dimensionality and semantic representation (Ma et al., 2024).
-Once processed, these embeddings are fed into the language model, specifically after the positional layer, to incorporate them within a sequential architecture.
-To enhance training efficiency and minimize computational cost, we freeze the language model’s parameters, making the MoE module the only trainable component in the architecture.
+A.3 Error Bound Analysis
+The approximations used in the proof rely on the assumption that $|x|$ and $|y|$ are large, which is common in high-dimensional hyperbolic embedding spaces.
+In practice, with embedding dimensions such as 50 (as used throughout our experiments), the norms tend to be sufficiently large for the approximation to hold effectively.
+To further substantiate this assumption, we calculated the error bound as follows:
+The exact Lorentz inner product is given by:
+$z=\sqrt{1+|x|^{2}}\sqrt{1+|y|^{2}}-|x||y|\cos\theta,$
+and is approximated as:
+$z_{\text{approx}}=|x||y|(1-\cos\theta).$
+Using a Taylor expansion, we have:
+$\sqrt{1+|x|^{2}}=|x|\left(1+\frac{1}{2|x|^{2}}-\frac{1}{8|x|^{4}}+\cdots\right).$
+Thus, the absolute error becomes:
+$|z-z_{\text{approx}}|=\frac{|y|}{2|x|}+\frac{|x|}{2|y|}.$
+The relative error in the Lorentz product is bounded by:
+$\epsilon_{z}\leq\frac{1}{2}\left(\frac{1}{|x|^{2}}+\frac{1}{|y|^{2}}\right)\frac{1}{1-\cos\theta}.$
+Additionally, the approximation error in $\sqrt{z^{2}-1}\approx z$ is given by:
+$\sqrt{z^{2}-1}=z\left(1-\frac{1}{2z^{2}}+\cdots\right),$
+which implies:
+Relative $\text{Relative error}\leq\frac{1}{2z^{2}}.$
+Combining these results, the overall error bound is:
+Relative Error in |$\text{Relative Error in }|\nabla d_{H}(x,y)|\leq\frac{1}{2|x|^{2}(1-\cos\theta)}+\frac{1}{2z^{2}}.$
+For instance, with $|x|=|y|=5$ and $\theta=\frac{\pi}{2}$, the total error is approximately $2.1\%$, demonstrating high approximation accuracy in our setting.
+Appendix B Experimental Settings
+In this section, we present details of our experimental settings, the distribution of head and tail items (Table 3), implementation details, and an explanation of the diversity metrics used.
+Notably, the high proportion of tail users (T80) in these datasets highlights the prevalence of long-tail distributions, reaffirming the assumption of a power-law distribution commonly observed in real-world data.
+B.1 Implementation Details
+We use an embedding dimension of 50 and a training batch size of 1024.
+The embeddings in hyperbolic space are optimized using Riemannian SGD (RSGD) [2, 25], while the MLP adapter is trained with the Adam optimizer [13].
+Training employs early stopping based on Recall@10, with a patience of 10 epochs.
+User and item profiles are generated using GPT-3.5-turbo and encoded with text-embedding-ada-002 [22] for the Amazon-Books and Yelp datasets, and BERT [8] for Google-Reviews.
+All experiments are conducted on NVIDIA A40/A100 GPUs with 40GB of memory.
+For the final experiments, we perform a grid search over commonly used hyperparameters: learning rate $\in\{1\mathrm{e}{-3},5\mathrm{e}{-3},1\mathrm{e}{-2},5\mathrm{e}{-2},1\mathrm{e}{-1}\}$, weight decay $\in\{1\mathrm{e}{-3},1\mathrm{e}{-2},1\mathrm{e}{-1}\}$, number of negative samples $\in\{10,20,50\}$, and alignment weight $\in\{1\mathrm{e}{-3},1\mathrm{e}{-2},1\mathrm{e}{-1}\}$.
+A full list of tuned hyperparameters is available in our attached GitHub repository.
+Table 3: Statistics of the experimental data.
+Dataset #User #Item #Interactions Density
+All H20(%) T80(%)
 […]
-B.1 Decoder Quality
-To assess the effectiveness of our decoder mechanism, we compare the embeddings from our model against those from Euclidean embeddings (LightGCN) and Hyperbolic embeddings (HICF) as inputs to the decoder.
-The results are shown in Table 5.
-We use BERTScore (Zhang et al., 2019) to evaluate the semantic alignment of generated sentences with the original profile.
-Results indicate that our HERec outperforms both Euclidean and hyperbolic baselines, demonstrating that our model’s embeddings support recommendation and retain user/item profile information.
-This finding underscores that our model captures collaborative information alongside recoverable language semantics: a capability not achieved by previous models.
-B.2 Decoder Case Study
-To vividly illustrate the semantic capacity of our embeddings, we present two case studies demonstrating the model’s ability to recover user and item profiles.
- Original user profile:
-Based on the user’s interactions and reviews, this user is likely to enjoy unique dining experiences with a variety of options, trendy atmosphere with good drink and food selections, premium quality frozen yogurt with a wide variety of toppings, creative burgers and hot dogs, delicious Thai cuisine, and authentic and spicy Chinese cuisine served family-style.
-Recovered user profile:
-Based on the user’s interactions, it is likely that they would enjoy businesses that offer a variety of cuisines, including Mexican, Italian, and Asian fusion.
-They appreciate good food, drinks, and a relaxed atmosphere.
-They also value quality and freshness in their food.
-In this example, the profile recovered by our decoder accurately captures the user’s preferences for Thai and Chinese cuisines, broadly summarized as an affinity for Asian fusion flavours.
-Additionally, our decoder identifies the user’s interest in diverse culinary experiences, indicating a propensity toward exploration.
- Original item profile:
-Ceol Irish Pub in Reno is a popular bar and pub known for its happy hour deals, great service, and friendly atmosphere.
-It is suitable for those who enjoy trivia nights, live music, and a wide selection of drinks, particularly whiskey.
-The pub has a welcoming environment, with bartenders who are attentive and quick to serve.
-Recovered item profile:
-Users who enjoy a variety of beer options, live music, and a relaxed atmosphere would enjoy Ceol Irish Pub.
-The alignment between the original and recovered item profiles shows that our decoder effectively captures key thematic elements of Ceol Irish Pub.
-The original profile emphasizes aspects such as happy hour, a wide selection of drinks, live music, and a welcoming atmosphere with attentive service.
-The recovered profile distills these characteristics, focusing on the pub’s appeal for users who enjoy diverse beer options, live music, and a relaxed setting.
-This alignment suggests that our decoder reliably summarizes key attributes, albeit with a degree of abstraction.
-Table 5: Experiments assessing decoder quality.
-Data Amazon-books Yelp
-BERTScore P R F1 P R F1
-Euclidean 0.3266 0.2854 0.3064 0.3511 0.2238 0.2872
-Hyperbolic 0.3498 0.3009 0.3257 0.3608 0.2398 0.3021
-Ours 0.3665 0.3167 0.3418 0.3716 0.2453 0.3081
-Appendix C Hierarchical Structure Case Study
-In this section, we present a case study examining the exploration-exploitation trade-off in the hyperbolic hierarchical structure.
+[…]
+[…]
+B.2 Diversity Metrics
+Distance Diversity (Div) quantifies the variation among recommended items, encouraging diversity within the recommendation:
+$Div(u)=\frac{\sum_{i\in R_{u}}\sum_{j\in R_{u},j\neq i}d(i,j)}{\text{Number of}(i,j)\text{ pairs}}$ (23)
+where $R_{u}$ is the recommendation list for user $u$, and $d(i,j)$ represents the distance between items $i$ and $j$.
+Number of $(i,j)$ pairs is calculated as $\binom{|R_{u}|}{2}=\frac{|R_{u}|(|R_{u}|-1)}{2}$.
+For fair comparison, we compute distances within the same representation space across models.
+Shannon Entropy (H) quantifies the unpredictability within the recommendations, with higher entropy signifying greater diversity:
+$H=-\sum_{i\in\text{set}(\sum_{u}R_{u})}p(i)\log p(i)$ (24)
+where $\sum_{u}R_{u}$ denotes all recommended items across users (including duplicates), and $p(i)=\text{count}(i)/\sum_{i\in\text{set}(\sum_{u}R_{u})}\text{count}(i)$ is the proportion of item $i$ among all recommended items.
+Expected Popularity Complement (EPC) assesses the extent of bias towards popular items by promoting niche recommendations:
+$EPC=1-\frac{1}{|\text{set}(\sum R_{u})|}\sum_{i\in\text{set}(\sum_{u}R_{u})}\frac{\text{pop}(i)}{\max_{i}(\text{pop}(i))}$ (25)
+where $N_{u}$ is the set of items user $u$ has interacted with, and $\text{pop}(i)=\text{count}(i)/\sum_{i\in\text{set}(\sum_{u}N_{u})}\text{count}(i)$ denotes item $i$’s popularity.
+Appendix C Supplement to Hyperbolic Hierarchical Clustering
+In this section, we provide details of the hyperbolic hierarchical clustering algorithm, together with complexity analysis and a detailed case study.
+C.1 Hyperbolic Hierarchical Clustering
+Algorithm 1 describes the hyperbolic hierarchical clustering procedure used to organize hyperbolic embeddings into a multi-layer structure.
+The algorithm takes as input the hyperbolic embeddings $X$ and a layer proportion parameter $k$, which controls the reduction rate of points per hierarchical layer.
+The maximum number of layers $L$ is set to $\log_{k}(|X|)$, ensuring that each subsequent layer contains approximately $1/k$ times the points of the previous layer.
+Starting from the bottom layer containing all points, the algorithm iteratively constructs higher layers in a bottom-up fashion.
+At each layer $l$, cluster centroids are randomly initialized in hyperbolic space, with the number of clusters equal to $|D_{l}|/k$, where $|D_{l}|$ is the number of points in the current layer.
+Clustering proceeds by repeatedly assigning each point to its nearest centroid (in terms of hyperbolic distance) and updating each centroid to the hyperbolic mean of its assigned points.
+This process continues until the centroids converge.
+Upon convergence, the centroids from the current layer serve as the points for the next upper layer.
+By repeating this process, the algorithm constructs a hierarchy of clusters, effectively capturing both the geometric structure and hierarchical relationships inherent in hyperbolic embeddings.
+Algorithm 1 Hyperbolic Hierarchical Clustering
+  Input: Hyperbolic embeddings $X$, layer proportion $k$
+  Set the maximum number of layers $L=\log_{k}(|X|)$
+  Initialize the points of bottom layer $D_{L}=X$
+  for $l=L$ to $1$ (bottom to top) do
+     Randomly initialize cluster centroids in the hyperbolic space $M_{l}=\{\mu_{l}\ |\ \mu_{l}\in\mathbb{H}\}$, $|M_{l}|=|D_{l}|/k$
+     repeat
+        for each point $d_{l,i}\in D_{l}$ do
+           Assign $d_{i}$ to the cluster with the nearest centroid:
+           $c_{l,i}=\arg\min_{\mu\in M_{l}}d_{\mathcal{H}}(d_{i},\mu)$
+        end for
+        for each cluster $c_{l,i}\in C_{l}$ do
+           Update the centroid $\mu_{l,i}$ as the hyperbolic mean $m_{\mathcal{H}}$ of all points assigned to cluster $c_{l,i}$:
+           $\mu_{l,i}=m_{\mathcal{H}}(\{d\ |\ d\in C_{l,i}\})$
+        end for
+     until Centroids $M_{l}$ do not change (convergence)
+     Initialize the points of the upper layer $D_{l-1}=M_{l}$
+  end for
+C.2 Complexity Analysis
+In this subsection, we analyze the complexity of constructing and applying the hierarchical representation structure.
+Let $N$ be the total number of users and items, and $k$ be the number of recommendations per user.
+Hierarchical Clustering.
+Building the hierarchy tree involves iterative k-means clustering in hyperbolic space.
+To optimize efficiency, we assign cluster nodes only to real node embeddings.
+First, we compute and store pairwise distances, requiring $O(N^{2}/2)$.
+Each k-means iteration takes $O(i)$, and with $O(\log N)$ levels, the overall complexity of constructing the hierarchy tree is $O(N^{2}/2+i\log N)$.
+Recommendation Retrieval & Replacement.
+Finding the top-$k$ recommendations requires $O(k\log N)$ using nearest neighbour search in hyperbolic space.
+Adjusting recommendations by retrieving ancestor clusters at layer $l$ requires $O(k)$ time, as it involves only tree traversal.
+The proposed hierarchical framework achieves efficient complexity scaling while maintaining flexibility in recommendation adjustments.
+The hierarchical clustering process, optimized for hyperbolic space, ensures a computationally feasible structure with $O(N^{2}/2+i\log N)$ complexity.
+The recommendation retrieval and replacement steps remain lightweight, running in $O(k\log N)$ and $O(k)$, respectively.
+This design enables HERec to dynamically balance exploration and exploitation while ensuring scalability for large-scale recommender systems.
+C.3 Hierarchical Structure Case Study
+In this subsection, we present a case study examining the exploration-exploitation trade-off in the hyperbolic hierarchical structure.
 We randomly select a user and identify items that share different layers of the lowest common ancestor (LCA) with the user in the hierarchy tree.
@@
 This analysis aligns seamlessly with the intended design of our hierarchical tree, further validating the effectiveness of our model in capturing and leveraging nuanced user-item relationships.
+Table 4: Comparison with feature-enhanced Euclidean baselines, with the best-performing results highlighted in bold.
+Data Amazon-books Yelp Google-reviews
+Metrics R@20 N@20 Div@20 H@20 EPC@20 R@20 N@20 Div@20 H@20 EPC@20 R@20 N@20 Div@20 H@20 EPC@20
+LightGCN+ 0.1466 0.0901 0.4349 10.3915 0.7802 0.1156 0.0733 0.2697 9.6903 0.7526 0.1438 0.0940 0.2707 10.8664 0.7801
+LightGCL+ 0.1545 0.0992 0.5064 11.9761 0.8728 0.1131 0.0739 0.2759 11.8617 0.8581 0.1443 0.0946 0.2744 12.4700 0.8776
+SimGCL+ 0.1494 0.0926 0.5163 11.4751 0.8342 0.1195 0.0752 0.3152 9.9952 0.7811 0.1504 0.0982 0.3030 11.2111 0.7953
+HCCF+ 0.1525 0.0940 0.4316 11.7984 0.8616 0.1170 0.0729 0.2884 11.6906 0.8577 0.1438 0.0941 0.2524 11.5433 0.8406
+Ours 0.1682 0.1069 0.5754 12.2543 0.8758 0.1224 0.0762 0.3961 12.3290 0.8773 0.1598 0.0992 0.3543 12.7690 0.8728
+Table 5: Standard deviation of key metrics across datasets over 10 independent runs.
+Dataset Recall@20 NDCG@20 DIV@20 H@20 EPC@20
+Amazon 0.0010 0.0004 0.0067 0.0035 0.0002
+Yelp 0.0009 0.0004 0.0091 0.0017 0.0014
+Google 0.0010 0.0003 0.0077 0.0008 0.0009
+Table 6: Direct message passing within hyperbolic space.
+Dataset Recall@20 NDCG@20 DIV@20 H@20 EPC@20
+Amazon 0.1721 0.1082 0.5760 12.211 0.8773
+Yelp 0.1205 0.0793 0.4005 12.321 0.8755
+Google 0.1611 0.0988 0.3532 12.766 0.8731
+Appendix D More Experimental Results
+In this section, we will show more experimental results to further analyze the flexibility and effectiveness of our HERec.
+D.1 Baselines with Feature Enhancement
+For further validation, we assessed the impact of feature integration on Euclidean baselines.
+By incorporating semantic embeddings into the Euclidean space—mirroring our approach in Section 4.2—we created enhanced variants (denoted by “+”).
+As reported in Table 4, these models generally achieve higher utility scores but do not consistently improve diversity.
+This reinforces our main findings: while feature integration benefits utility in Euclidean frameworks, it falls short for diversity, highlighting the unique strengths of hyperbolic alignment detailed in Section 4.4.
+D.2 Statistical Significance Testing
+To ensure the robustness and reliability of our results, we conducted 10 independent runs for each experimental setting and reported the standard deviation for all key metrics.
+Results are shown in Table 5.
+Across all datasets and evaluation metrics, we observed consistently low standard deviations, indicating high stability and reproducibility of our findings.
+This statistical rigor further substantiates the significance of our experimental conclusions.
+D.3 Alternative Hyperbolic Message Passing
+We also explored an alternative message passing strategy by replacing the original hyperbolic graph collaborative filtering mechanism with direct message passing in hyperbolic space, thereby eliminating the need for tangent space projections.
+This variant achieves performance comparable to our primary method, results can be found in Table 6.
+These findings suggest that direct message passing in hyperbolic space is not only a viable alternative but could also serve as a complementary component within our framework, potentially offering additional flexibility for future extensions.
+Appendix E Profile Decoder
+System Instruction: Describe what types of items this user is likely to enjoy./ Describe what types of users would enjoy this item.
+Input Prompt: <USER_EMBED>/<ITEM_EMBED> Output: The user would prefer items which …/ The item would be preferred by users who …
+Figure 5: A depiction of model prompt instruction.
+To further evaluate the effectiveness of hyperbolic alignment and the expressiveness of our learned hyperbolic embeddings, we introduce a profile decoder designed to reconstruct the original profile from these embeddings, thereby broadening their application beyond recommendation tasks.
+Our method begins by mapping the hyperbolic embeddings back into Euclidean space to facilitate alignment with traditional network layers.
+These Euclidean-mapped embeddings are then processed through a mixture of experts (MoE) module.
+The MoE serves to refine the embeddings, ensuring consistency in both dimensionality and semantic representation [20].
+Once processed, these embeddings are fed into the language model, specifically after the positional layer, to incorporate them within a sequential architecture.
+To enhance training efficiency and minimize computational cost, we freeze the language model’s parameters, making the MoE module the only trainable component in the architecture.
+[…]
+E.1 Decoder Quality
+To assess the effectiveness of our decoder mechanism, we compare the embeddings from our model against those from Euclidean embeddings (LightGCN) and Hyperbolic embeddings (HICF) as inputs to the decoder.
+The results are shown in Table 7.
+We use BERTScore [44] to evaluate the semantic alignment of generated sentences with the original profile.
+Results indicate that our HERec outperforms both Euclidean and hyperbolic baselines, demonstrating that our model’s embeddings support recommendation and retain user/item profile information.
+This finding underscores that our model captures collaborative information alongside recoverable language semantics: a capability not achieved by previous models.
+E.2 Decoder Case Study
+To vividly illustrate the semantic capacity of our embeddings, we present two case studies demonstrating the model’s ability to recover user and item profiles.
+Original user profile:
+Based on the user’s interactions and reviews, this user is likely to enjoy unique dining experiences with a variety of options, trendy atmosphere with good drink and food selections, premium quality frozen yogurt with a wide variety of toppings, creative burgers and hot dogs, delicious Thai cuisine, and authentic and spicy Chinese cuisine served family-style.
+Recovered user profile:
+Based on the user’s interactions, it is likely that they would enjoy businesses that offer a variety of cuisines, including Mexican, Italian, and Asian fusion.
+They appreciate good food, drinks, and a relaxed atmosphere.
+They also value quality and freshness in their food.
+In this example, the profile recovered by our decoder accurately captures the user’s preferences for Thai and Chinese cuisines, broadly summarized as an affinity for Asian fusion flavours.
+Additionally, our decoder identifies the user’s interest in diverse culinary experiences, indicating a propensity toward exploration.
+Original item profile: Ceol Irish Pub in Reno is a popular bar and pub known for its happy hour deals, great service, and friendly atmosphere.
+It is suitable for those who enjoy trivia nights, live music, and a wide selection of drinks, particularly whiskey.
+The pub has a welcoming environment, with bartenders who are attentive and quick to serve.
+Recovered item profile:
+Users who enjoy a variety of beer options, live music, and a relaxed atmosphere would enjoy Ceol Irish Pub.
+The alignment between the original and recovered item profiles shows that our decoder effectively captures key thematic elements of Ceol Irish Pub.
+The original profile emphasizes aspects such as happy hour, a wide selection of drinks, live music, and a welcoming atmosphere with attentive service.
+The recovered profile distills these characteristics, focusing on the pub’s appeal for users who enjoy diverse beer options, live music, and a relaxed setting.
+This alignment suggests that our decoder reliably summarizes key attributes, albeit with a degree of abstraction.
+Table 7: Experiments assessing decoder quality.
+Data Amazon-books Yelp
+BERTScore P R F1 P R F1
+Euclidean 0.3266 0.2854 0.3064 0.3511 0.2238 0.2872
+Hyperbolic 0.3498 0.3009 0.3257 0.3608 0.2398 0.3021
+Ours 0.3665 0.3167 0.3418 0.3716 0.2453 0.3081